### v1

# A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War

Abstract
We introduce dnipro, a novel longitudinal corpus of 246K news articles documenting the Russo-Ukrainian war from Feb 2022 to Aug 2024, spanning eleven media outlets across five nation states (Russia, Ukraine, U.S., U.K., and China) and three languages (English, Russian, and Mandarin Chinese). This multilingual resource features consistent and comprehensive metadata, and multiple types of annotation with rigorous human evaluations for downstream tasks relevant to systematic transnational analyses of contentious wartime discourse. dnipro’s distinctive value lies in its inclusion of competing geopolitical perspectives, making it uniquely suited for studying narrative divergence, media framing, and information warfare. To demonstrate its utility, we include use case experiments using stance detection, sentiment analysis, topical framing, and contradiction analysis of major conflict events within the larger war. Our explorations reveal how outlets construct competing realities, with coverage exhibiting polarized interpretations that reflect geopolitical interests. Beyond supporting computational journalism research, dnipro provides a foundational resource for understanding how conflicting narratives emerge and evolve across global information ecosystems.
A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War
Dikshya Mohanty, Taisiia Sabadyn, Jelwin Rodrigues, Chenlu Wang, Abhishek Kalugade, Ritwik Banerjee Stony Brook University Correspondence: {dimohanty, jelrodrigues, rbanerjee}@cs.stonybrook.edu
1 Introduction
The Russo-Ukrainian war, which escalated on February 24, 2022 with the declaration of a "special military operation" by the Russian government, [^1] Within 48 hours, media in Russia were banned from describing the developments as “war” AFP (2022). has significantly impacted the global geopolitical landscape Derviş (2023); Silvertsen (2024). The war has redrawn traditional alliances, tested the resolve of international institutions, and created new geopolitical fault lines. Despite initial expectations of a swift Russian victory, Ukraine’s resistance has resulted in a protracted war with substantial loss of life and human capital Watling et al. (2023); Pancevski (2024). This conflict has also been characterized by a quieter battle for narrative control, with news coverage from several nations playing a pivotal role in shaping global public opinion and influencing the war’s trajectory Bradshaw et al. (2024). These sources reflect their national priorities, strategic perspectives, and the perceptions they intend to create among information consumers Johais and Meis (2024). Yet, despite the centrality of media and narrative framing, there remains a lack of comprehensive datasets for longitudinal analysis of how nations have framed the war for their domestic and international audiences.
We introduce a collection of Diverse Narratives and International Perspectives on the Russo-Ukrainian Offensive (dnipro), a comprehensive longitudinal corpus of news coverage from February 2022 to August 2024222An expanded version is being developed, which will include news coverage up to February 2025 and provide a longitudinal 3-year corpus of this protracted war.—the longest time span of any multinational or multilingual news corpus on the Russia-Ukraine war. The corpus encompasses eleven influential media outlets and news agencies aligned with five primary geopolitical actors with strategic interests in the conflict. Following methodological approaches established by Nygren et al. (2016) and Roman et al. (2017), the sources include media from direct participants—Russia and Ukraine—and influential global powers with economic, diplomatic, or military involvement: USA, UK, and China. The inclusion of state-affiliated media from Russia and Ukraine capture opposing conflict narratives, while those from China, the USA, and the UK represent the narratives projected by nations with significant stakes in the conflict’s outcome, and significant ability to influence that outcome Zhang et al. (2016); Thussu et al. (2018).
dnipro offers rich annotations for entity-level stance, sentiment, and topical framing, enabling event tracking and other longitudinal analyses across geopolitical perspectives. The dataset does not attempt to verify the accuracy of any claims; rather, it provides a resource for studying how wartime narratives are constructed and disseminated by major geopolitical actors. [^3] This aligns with established methodologies in discourse analysis Entman (1993); Fairclough (2013) examining how language and framing construct social reality. Further, Gamson and Modigliani (1989) showed how this view offers a powerful tool to understand the competition that shapes public understanding of contentious events. The corpus is uniquely suited to such studies, as it comprises 246,229 articles across 11 media sources and 3 languages, spanning 5 diverse nations, over a period of 31 months. Our key contributions include:
  • (1)
    a longitudinal multinational and multilingual corpus of news coverage spanning the first 30 months of the Russia-Ukraine war;
  • (2)
    annotations supporting comparative analyses of named entities, stance, and framing; and
  • (3)
    the inclusion of diverse media sources enabling cross-cultural examination of geopolitical narratives.
Next, we place dnipro within the landscape of existing conflict-related corpora and highlight its distinct contributions to the field (§ 2), and describe how our design and construction choices (§ 3) make it a powerful resource for large-scale analyses of wartime narratives. We employ dnipro in several applications, using state-of-the-art language understanding systems as well as additional annotations (§ 4 and § 5), demonstrating its utility.
2 Related Work
Research on wartime media coverage relies on datasets capturing narratives across national and institutional contexts, where competing claims and propaganda techniques pose unique challenges. Such resources are vital for training models to analyze geopolitical discourse and evaluating their ability to identify narrative strategies. As a reflection of diverse goals, existing resources vary in scope, modality, and annotations.
Twitter & Reddit-based datasets are among the largest. Chen and Ferrara (2023) aggregate over 570 million tweets in English, Russian, and Ukrainian to study information propagation and influence operations, while Zhu et al. (2022) provide 8 million posts correlating post categories with pro-Ukrainian stance. Shevtsov et al. (2022) introduce 57.3 million English tweets annotated for sentiment and hate speech, and the Invasion@Ukraine corpus provides 8.7 million English tweets for content moderation and conspiracy narrative analysis Susanne Pohl et al. (2023). The VoynaSlov collection Park et al. (2022) bears similarities to our work through sentence-level media framing labels, though they focus solely on Russian media.
Alyukov et al. (2023) introduce one of the first cross-platform social media datasets, integrating 1.7 million posts to enable propaganda and public opinion analyses. Telegram-specific corpora include 4.1 million posts with pro/anti-Kremlin channel tags Bawa et al. (2025) and 2.3 million posts with binary tags for human rights violations Nemkova et al. (2023).
Hakimov and Cheema (2024) introduce a corpus of 1.5 million tweets in 60 languages with automated sentiment, stance, and named entity annotations, offering the widest multilingual social media dataset. The RUWA corpus Khairova et al. (2024) provides 16,500 news articles using semantic similarity for misinformation detection, but excludes the overarching war theme and curates only specific event-based articles. dnipro supersedes it by orders of magnitude in coverage, topics, multilinguality, temporality, and size. Ibrahim et al. (2025) offer longitudinal media analysis of two years of news data, but from nations with auxiliary involvement.
Annotation practices differ widely: manual schemata for framing Park et al. (2022), channel stance Bawa et al. (2025), and human rights violations Nemkova et al. (2023); automated annotations for large-scale labeling Susanne Pohl et al. (2023); Hakimov and Cheema (2024). Most resources rely on social media identifiers that degrade quickly, impeding reproduction. Few provide longitudinally contiguous, multilingual, news-based content annotated for multiple tasks.
dnipro extends existing corpora by focusing on transnational and multilingual news coverage – underrepresented in prior social media-centered work – enabling comparative analyses of state-aligned narratives across geopolitical actors. It reflects news coverage as disseminated, without alteration, serving as a neutral resource for empirical studies. Our design choices spanning source selection, temporal coverage, and structured annotations make dnipro a powerful resource for large-scale automated analysis of wartime narratives (§ 5).
3 The dnipro Corpus
Our dataset is constructed through a systematic multi-stage pipeline targeting news outlets and state-affiliated agencies selected on the basis of three criteria: (1) global influence and prominence within national media ecosystems, (2) sustained coverage of the Russo-Ukrainian war, and (3) discernible editorial positioning ranging from independent journalism to state-aligned messaging. To the extent possible, access to full article archives was obtained through public RSS feeds, institutional subscriptions, or source-specific data collection pipelines. However, not all media outlets provided consistent or unrestricted access to their archives, and some selections reflect practical constraints in addition to topical relevance.
3.1 Data Acquisition
dnipro is a corpus spanning 31 months of daily data collection, from February 1, 2022 to August 31, 2024. To identify articles with substantive coverage of the Russo-Ukrainian conflict, we employ a tiered strategy. For outlets with dedicated sections or labels related to the war, we collect all articles published under those categories. Where available, we also extract articles tagged with conflict-specific metadata (e.g., “Ukraine conflict”, “special military operation”, “Russia-Ukraine war”). For sources lacking such taxonomies, we apply keyword-based filtering to isolate relevant content. However, some sources impose significant restrictions on retrospective API-based keyword searches. In these cases, we obtain the older articles directly through web browsers, operating within the scope of respective terms of service and copyright restrictions. To ensure quality control, we manually validate stratified random samples across sources, languages, and time periods throughout the acquisition process.
Total no. of articles 246,229
Title length (avg.) 11.7 words
Article length (avg.) 332.5 words
No. articles/day (avg.) 261.1
No. of sources 11
Languages covered N/A
Temporal span Feb 1, 2022 - Aug 31, 2024
No. of non-English articles 61,637 (25.03%)
Geopolitical entities Russia, Ukraine, China, USA, UK
After translation (of non-English sources to English).
Table 1: Key descriptive statistics of dnipro.
English-language “Western” media from the US and the United Kingdom, namely The Financial Times (FT) and Cable News Network (CNN), were included as part of the Western bloc, reflecting the perspectives of countries with significant political, economic, and military involvement in the conflict. These outlets serve domestic and international audiences and are central to understanding how Western-aligned narratives are constructed and disseminated. No additional filtering of articles is done by search queries as these media houses offer dedicated sections for content related to the Russo-Ukrainian war. We collect a total of 5,760 and 4,218 articles from CNN and FT, respectively.
To ensure adequate representation of Chinese state-aligned perspectives, we collect articles from three leading outlets: Xinhua, Global Times, and China Daily. Xinhua publishes exclusively in Mandarin, whereas Global Times and China Daily provide parallel English-language coverage. Articles are collected from Xinhua using a bilingual keyword list, designed to minimize the loss of coverage while retaining topical precision. Mandarin Chinese queries comprised “乌克兰” (Ukraine), “俄罗斯” (Russia) and “俄乌” (Russia and Ukraine); English queries included “Russia–Ukraine”, “Russia”, “Ukraine”, and “Ukrainian”. These queries yield in 11,152 and 10,206 articles in Mandarin Chinese and English, respectively.
[…]
The Ukrainian perspective is represented by articles from Censor.net, The Kyiv Independent (KI), and European Pravda (EP), all of which provide English-language versions. While KI and EP shape Ukraine’s global image, Censor influences the domestic audience. Their reporting styles vary—from Censor’s rapid updates to EP’s in-depth opinions and interviews. As sources targeting the global Anglosphere, KI and EP offer sections dedicated to the Russo-Ukrainian War: “War” and “War with Russia”, respectively. Censor’s news reporting approach largely covers the war, evidenced by the majority fraction of Ukrainian war-related tags in their “News” section. As such, no keyword search queries are applied to any of the Ukrainian sources. The collection from these courses amount to 19,706 articles from Kyiv Independent, 2,212 from European Pravda, and 31,748 from Censor.
The Russian media sources in our dataset comprise English articles from TASS and Sputnik, and Russian-language articles from Izvestia. No keyword search queries are required, as all articles are collected from the dedicated section “Special Operation of Russia in Ukraine”. We collect 60,002 and 53,139 articles from TASS and Sputnik respectively, and 50,485 articles from Izvestia.
article_id A unique UUID v4 identifier for the record.
url The persistent URL from which the article was sourced.
publication_date The article’s publication date in ISO 8601 format (yyyy-mm-dd).
source_name The normalized name of the news outlet (e.g., Xinhua).
source_country The country of origin for the news outlet, in ISO 3166-1 alpha-3.
language The article language, in ISO 639-1.
title The article’s title (i.e., headline).
article_length The number of words in the cleaned article text (post processing).
collection_date The date (in ISO 8601 format) when the URL was accessed.
Table 2: Core metadata fields for each article.
3.2 Metadata Schema and Statistics
The core of dnipro is a collection of rich, consistent metadata records for 246,229 news articles. Its key descriptive statistics are shown in Table 1. Each record is described by a unified schema (described in Table 2) designed to facilitate large-scale longitudinal and comparative analyses.
The overall composition of the natural language data—both in terms of the number of news articles and their sizes—across individual news sources and aggregated by geopolitical entities, is presented in Figure 1.
3.3 Data Records
dnipro is publicly available on Zenodo under a CC-BY-NC 4.0 license. [^4] The data, along with all processing code and annotation prompts, will be released upon acceptance. It is structured for ease of use and distributed as a single, compressed tar.bz2 archive containing a primary metadata file, annotation files, and a documentation file. [^5] We do not directly include the texts due to copyright restrictions. The provided URLs serve as persistent identifiers enabling reproduction of dnipro, like most social media datasets. If an article becomes unavailable, the collection_date and url enable retrieval via the Internet Archive’s WayBack Machine.
  1. 1.
    metadata.parquet contains the complete set of 246,229 metadata records as described in § 3.2. The Parquet format was chosen for its efficient compression and columnar storage, which enables fast querying and analysis using popular data science tools.
  2. 2.
    named-entities.parquet contains the named entity annotations for the entire corpus. Each record includes an article_id (see Table 2) and the list of named entities in that article.
  3. 3.
    sentiments.parquet provides sentiment scores for select entities and events across multiple news sources. For politically important and frequently occurring entities (e.g., Ukraine, Russia, Zelensky, Kherson), these scores are aggregated monthly. For entities tied to key events within this war (e.g., Bucha, Kursk, Zaporizhzhia), they are aggregated daily.
  4. 4.
    topical-frames.parquet provides one row per article_id with its corresponding topical frames anchored in the nine-category schema described in Table 3.
  5. 5.
    README.md is a comprehensive description of the dataset, the schemata, the license terms, and a code of conduct for ethical use.
We do not specify train/test/dev splits for dnipro. The corpus spans time, languages, and multiple sources; and supports diverse study designs (e.g., temporal evolution or development of an entity/topic/event; media source comparisons; transnational or cross-lingual reporting of events; or other types of aggregation). A preset split would constrain the diversity of possible applications. We encourage task-appropriate partitions instead.
4 Data Annotations
We now describe the enrichment of raw article metadata with annotations. Since copyright restrictions prevent the direct release of texts, we generate content-derived labels that form the dataset’s core analytical value. The process involves:
  1. 1.
    metadata and text normalization;
  2. 2.
    translation of non-English text to English;
  3. 3.
    labeling named entities;
  4. 4.
    sentiment analysis;
  5. 5.
    stance analysis;
  6. 6.
    topical framing; and
  7. 7.
    rigorous quality control with human evaluation.
4.1 Text Processing for Annotation
To prepare the corpus for annotation and analyses, we apply standardized processing steps to all data, across sources and languages. This ensures consistency across media sources with varying conventions in structure, formatting, and metadata.
Publication dates are parsed and normalized to ISO 8601 standard (yyyy-mm-dd), and multiple metadata fields—e.g., “publication date”, “published on”—are mapped to metadata fields in our unified schema (see Table 2). Text content is stripped of non-informative boilerplate elements like location or date prefixes commonly found in news articles (e.g., “KYIV, March 3 –”), inline footnotes, bylines, and repeated source disclaimers. Duplicate articles are removed, to avoid redundancy in print syndication and multilingual reprints. To support downstream linguistic analyses, the text is then converted to UTF-8 encoding, segmented into paragraphs, and split into sentences.
Frame Definition
economy Macroeconomic issues and markets: sanctions/trade, inflation, growth, jobs, taxes, energy/commodity prices.
identity Group identity and belonging: nationality, ethnicity, religion, language, or gender; discrimination or identity politics.
morality Explicit ethical/value framing: moral duty, fairness, dignity, humanitarian arguments.
legal Courts and legal processes: rulings, lawsuits, prosecutions, designations, warrants, legal rights or entitlements.
policy Concrete government outlook or outreach, rules, bills, decrees, program design and details (what a policy does).
politics elections, party/elite competition, horse-race strategy, reshuffles, coalition dynamics.
public opinion Polls, surveys, attitudes and protests as expressions of public sentiment.
security War/defense, violence/crime, policing, border enforcement, terrorism, cyberattacks.
welfare Benefits and human services: healthcare, pensions, housing aid, unemployment, assistance to refugees or internally displaced people.
Table 3: Topical frame definitions in dnipro: a significant abridgment of the Refugees and Migration Framing Vocabulary Yu and Fliethmann (2022).
4.2 Machine Translation
A quarter of dnipro consists of Mandarin Chinese and Russian texts, (zh and ru, respectively, in ISO 639-1). We translate these into English (en) to facilitate downstream analyses. We choose three models based on recent comparative studies on neural machine translation by Smirnov et al. (2022) and Alemayehu et al. (2024): OPUS-MT Tiedemann et al. (2023), M2M100 Fan et al. (2020), and NLLB-200-Distilled Team (2022). Since there were no readily available explicit benchmarks for zh-to-en or ru-to-en translations, we select a random set of ten articles from Xinhua, evaluated by two native speakers of Mandarin Chinese for semantic fidelity and fluency. The evaluators worked independently to score the output of these models on 1-5 Likert scale for fidelity and fluency, and M2M100-418M received the highest scores from both evaluators on both criteria. So, we employ this model for all zh-to-en translations. Russian articles were similarly translated using the ru-to-en path of M2M100-418M.
4.3 Annotation Methodology
To enable fine-grained analysis of key actors and locations, we perform Named Entity Recognition (NER) across the entire corpus (after translation to English). This annotation is carried out using the spaCy library Honnibal et al. (2020) and its pretrained model en_core_web_sm, identifying entities and their types according to the OntoNotes 5.0 taxonomy Weischedel et al. (2013). This process results in the identification of 85,155 distinct named entity mentions, and provides a foundational layer for tasks such as tracking the prominence of specific actors over time, temporal studies of media focus across entities and events, and entity-specific sentiment and stance analyses.
Event (Date range) # articles
Bucha Massacre (Apr 3–9, 2022) 273
Zaporizhzhia Nuclear Crisis (Mar 3–9, 2022) 219
Mariupol Hospital Attack (Mar 9–15, 2022) 138
Kursk Offensive (Aug 6–12, 2024) 180
Table 4: The four conflict events analyzed in our use case experiments, selected based on global coverage and the potential to generate narrative divergence.
To capture the sentiment of some of the most prominent named entities, we provide sentence-level sentiment annotation using a pretrained model: sentiment-roberta-large-english Hartmann et al. (2023). For each article, sentences containing such entities (e.g., Kursk, Bucha, Kherson, Zelensky, Ukraine, Russia) are identified using spaCy, and passed to this classifier. The model outputs are then mapped to continuous sentiment scores, which are then aggregated on a daily and monthly basis at both the source and country level using source–country relations. These annotations thus provide time series data revealing sentiment trends for each entity.
The dataset includes stance annotations obtained by few-shot instruction prompting with Phi-4 14B model  Microsoft (2024), identifying the stance towards a specified target. The input comprises a sentence and the immediate context in its news article. As part of the input, we provide a task instruction with a few examples, along with (i) a passage for context resolution, (ii) a focal sentence, and (iii) the target. The model is prompted to output a compact, formatted response:
Target: <the identified target>
Stance: favor|against|none
Rationale: <a short justification grounded in the given text>
For topical framing, we again use few-shot prompting with Phi-4 14B model, where the input is an instruction describing the task, along with label definitions and 1-2 examples per label. Then, for a given instance – an article’s title, subtitle (if present) and content – we prompt the model to generate a json object containing the list of applicable topical frames, and associated confidence scores.
4.4 Human Evaluation
To evaluate the reliability of our automated annotations, we conduct human evaluations of the stance and topical frame annotations.
Frame PABAK Confidence
Human/Human Human/LLM Human LLM
Economy 0.70 (84.8%) 0.736 0.84 0.93
Identity 0.74 (86.8%) 0.783 0.79 0.92
Legal 0.85 (92.4%) 0.783 0.74 0.91
Morality 0.75 (87.3%) 0.750 0.79 0.91
Policy 0.55 (77.7%) 0.526 0.78 0.89
Politics 0.48 (74.1%) 0.425 0.79 0.92
Public Opinion 0.87 (93.4%) 0.865 0.80 0.89
Security 0.69 (84.3%) 0.587 0.85 0.93
Welfare 0.81 (90.4%) 0.783 0.80 0.87
Average 0.71 (85.7%) 0.693 0.798 0.909
Table 5: Inter-annotator agreement and confidence by topical frame. Pabak columns report human inter-annotator agreement (percent agreement in parentheses) and human–model agreement. Confidence columns report average confidence scores from human annotators and LLM (Phi-4).
Topical Frames:
Two expert annotators independently label a stratified random sample of 200 articles, providing nine binary labels with corresponding [0,1] confidence scores for each article. Significant class imbalance is a common challenge in multi-label annotation, and the inter-annotator agreement must consider the high base rate agreement where annotators agree on a label being absent from an article. The average agreement on label absence, over all nine topical frames, accounts for 65.82% of our evaluation sample, with the positive class (for any label) representing a minority of instances. Due to this prevalence bias, traditional measures like Cohen’s or Fleiss’ kappa are overly pessimistic reliability measures for our data. For a more accurate assessment that accounts for this imbalance, we report the prevalence-adjusted bias-adjusted kappapabak Byrt et al. (1993) – which is especially crucial for evaluation of consensus on the rarer labels.
The inter-annotator agreement results, measured by pabak, are presented in Table 5 and demonstrate substantial agreement on average (0.71, with 85.7% agreement). Three topical frame categories — Welfare, Public Opinion, and Legal — reveal near-perfect agreement (>> 0.8). Overall, seven of the nine labels achieve pabak scores $\geq$ 0.7, indicating a consistent and shared understanding of the annotation framework. The two remaining classes, Policy and Politics, have moderate agreement, as they represent more abstract and often overlapping conceptual domains. We also examine the correlation between the average confidence scores reported by the human experts and those reported by Phi-4. The framing categories are complex and sometimes overlapping, resulting in a weak positive correlation $\rho$ = 0.37.
Implications for dataset use. The strong agreement validates the overall reliability of our annotation schema. However, we recommend that users consider the tiered nature of this reliability when designing their experiments:
  • The seven frames with pabak > 0.69 are suitable for fine-grained analysis, including temporal tracking and cross-source comparison.
  • The Policy and Politics labels yield lower agreement due to their inherent conceptual overlap. We recommend these labels be used for high-level topic modeling or as supplementary features in larger models, with the understanding that they represent broader and possibly intersecting, categories of discourse.
  • The human-annotated sample of 200 documents, while modest in size, provides a high-quality validated benchmark. It was strategically selected through stratified sampling to ensure longitudinal representation across sources and languages, making it suitable for evaluating the performance of topical frame classification systems on real-world data on contentious wartime narratives.
Annotator-pair Agreement (%) Cohen’s $\kappa$
Human annotators 86.00% 0.770
Annotator1 and phi-4 72.68% 0.618
Annotator2 and phi-4 70.00% 0.580
$\bullet\;\;$ Mean agreement: 76.23%
$\bullet\;\;$ Fleiss’ $\kappa$ across all three annotators: 0.643
$\bullet\;\;$ Mean pairwise Cohen’s $\kappa$: 0.656 (range: 0.58-0.77)
Table 6: Inter-annotator agreements for stance labels, measured by percentage of agreement, pairwise Cohen’s $\kappa$, and Fleiss’ $\kappa$.
Stance Annotations:
The stance annotation task identifies whether a sentence (with context) expressed a stance (favor or against) towards any entity, or if no stance was present (none). Two human annotators with native proficiency in English annotated a sample of 150 instances. We evaluate their inter-annotator agreement, and then also with phi-4 as the LLM annotator. Table 6 presents the pairwise and overall inter-annotator agreements, which indicate strong human inter-annotator agreement ($\kappa$ = 0.77). The LLM annotator achieved moderate agreement with both ($\kappa$ = 0.62 and 0.58), showing consistent but less-than-human reliability. Fleiss’ $\kappa$ = 0.643, however, suggests that the stance annotations are sufficiently reliable for large-scale analysis, despite the gap between human and automated annotation quality for this task.
Figure 2: Topical framing by source, with % share of articles for each label (confidence >> 0.5). Source abbr.: Financial Times (FT), Kyiv Post (KP), Sputnik (SP), European Pravda (EP), Global Times (GT), China Daily (CD), Xinhua (XH). Labels: (Sec)urity, (Pol)itics, (Econ)omy, (Welf)are, (Ident)ity, (Moral)ity, and (Pub)lic (Op)inion.
5 Use Case Experiments
To demonstrate the analytical potential of dnipro, we present several studies that leverage its unique multi-perspective structure. These experiments are intended as use case demonstrations of the dataset’s capacity to support diverse research questions in computational journalism, political science, and discourse analysis. Each application highlights specific affordances of the corpus while suggesting broader research directions.
Thematic Analysis of Conflict Events:
Our analyses focus on four major conflict events that received substantial global coverage and generated clear geopolitical divisions, inviting transnational comparisons (see Table 4). For each event, we filter articles using specific location identifiers (e.g., Bucha) to ensure event-specific analysis.
To examine how media outlets prioritize different themes when covering conflict events, we analyze the topical content and emphasis patterns across sources. We identify which concepts, actors, and subtopics receive attention in coverage of major events, revealing how outlets differentially emphasize aspects such as historical context, humanitarian concerns, or institutional responses.
Using GPT-3.5 Ye et al. (2023) with structured prompts, we tag articles from weekly time windows surrounding each major event to identify key concepts, actors, and subtopics. For the events described in Table 4, we obtain 810 articles, with 10.5 such tags (avg.) per article. We aggregate these tags by event and source country to find the distinct thematic emphasis patterns across media outlets.
Russian and Chinese outlets consistently emphasize institutional terminology (special military operation), security concerns (nuclear weapons), and counter-narratives (media manipulation, ceasefire), reflecting strategic adversarial framing. Ukrainian sources predominantly highlight humanitarian dimensions (civilian casualties) with frequent references to specific conflict locations. American and British media balance humanitarian topics with institutional responses (sanctions, international criminal court) and policy discussions (no fly zone).
These patterns demonstrate how dnipro captures differences in thematic prioritization across geopolitical actors even when covering the same underlying events.
Russian Media Ukrainian Media
Russia had testimony from residents of Bucha and Mariupol, who reported brutal crimes, such as torture, committed by Ukrainian nationalists. (TASS) All eyewitnesses noted the systematic nature of the killings in Bucha by the Russian army. (Censor)
Russian forces jointly with border guards repelled these attacks and hit enemy positions near the border and in the Sumy Region. (TASS) Russian forces attacked six border areas and settlements of Sumy Oblast on June 7, firing 14 times and causing at least 54 explosions, the Sumy Oblast Military Administration reported. (Kyiv Independent)
Table 7: Contradictory claims from opposing media outlets, highlighting how conflicting narratives are constructed not through direct refutation, but via divergent attribution, framing, and selection biases.
Topical Framing:
We conduct a corpus-wide analysis using our frame definitions (Table 3). Article-level topical frames provide a comparable map of what each media outlet emphasizes, revealing systematic priorities that contextualize downstream analyses such as stance, sentiment, or discourse. Articles are annotated with phi-4 using a few-shot instruction prompt, producing multi-label assignments along with confidence scores.
Figure 2 shows that Security dominates across all sources, with the highest shares in Ukrainian media, followed by Russian state media and Western sources like CNN. The second-most prevalent category is Politics, elevated in Russian and Chinese sources. It is substantially present in other sources too. Economy is most visible in Chinese media and, naturally, The Financial Times, while Policy appears steadily across sources, with higher emphasis in Chinese and British coverage. Relative to others, Russian media leans more on Legal topics. Notably, discussions of Morality are extremely rare despite the focal topic of this dataset being an extended war, indicating that operational, strategic, and material considerations far outweigh explicit ethical rhetoric in war coverage.
Figure 3: Daily average of stance scores towards Ukraine, grouped by media sources from five nations. The first 15 days of the war (left), contrasted with the last 15 days collected for dnipro (right).
Stance analysis:
Based on our stance annotations in § 4.3, we provide a comparison of the stance towards Ukraine between the first fortnight of the war and the last 15 days of dnipro data. This reveals a striking temporal shift across geopolitical lines. In Figure 3, we see that Chinese media was slightly pro-Ukrainian. By August 2024, however, this shifted to mild opposition. Russian coverage is consistently negative, but there is a drastic negative shift. The US media stance becomes even more supportive, while UK news begins to exhibit greater variability. This temporal comparison demonstrates how dnipro can provide analytical insights into reporting patterns. In this use case, it illustrates the polarization of geopolitical alignments.
Divergent Narratives:
To investigate how media outlets construct and contest narratives around specific flash points in the Russo-Ukrainian war, we analyze claims to identify conflicting representations of key events. Building on prior work in conflict journalism Entman (2008); Thussu et al. (2018), this analysis unearths how opposing perspectives are embedded in individual claims.
Using weekly data subsets for each event, we extract sentence-level claims and categorize them into six types adapted from prior factuality and attribution typologies (Kotonya and Toni, 2020; Spinde et al., 2021):
  • (1)
    direct statements of observable or confirmed events (factual);
  • (2)
    claims attributed to specific actors (attributed);
  • (3)
    explicit denials or rebuttals (disputed);
  • (4)
    indirect accounts from third-party media (reported);
  • (5)
    official statements of intent or directive (command/policy); and
  • (6)
    interpretive or evaluative judgments by experts or institutions (analytical/assessment).
With GPT-3.5 plus human-in-the-loop validation, we label claims via structured prompts, and group semantically related claims through agglomerative clustering on Sentence Transformer representations Reimers and Gurevych (2019).
This approach reveals distinct narrative patterns across geopolitical boundaries: Russian sources might characterize an airstrike as precision targeting of military assets (factual/attributed), while American, British, or Ukrainian outlets describe the same event as civilian infrastructure targeting, often citing eyewitnesses (disputed/reported). Analyzing claim pairs shows entailment occurring primarily within aligned media ecosystems: Ukrainian and Western sources reinforce each other through overlapping factual, reported, or command/policy claims, while Russian media maintains internal consistency through attributed and analytical claims citing official statements. Contradictions predominantly emerge across geopolitical divides as competing interpretations rather than direct disputes. Outlets rarely offer explicit rebuttals, instead employing different topical framing, omission, or counter-attribution strategies.
Table 7 presents contradictory claim pairs from Russian and Ukrainian media. While referring to the same events, the claims construct incompatible interpretations of responsibility and action through attribution and selective emphasis.
6 Conclusion
The dnipro corpus represents a significant advancement in resources for computational analysis of media narratives during international conflicts. By integrating diverse geopolitical perspectives, consistent translations, and structured metadata across 246K articles, this dataset enables systematic study of narrative divergence that previous resources could not support. Our exploratory applications demonstrate dnipro’s utility across multiple analytical dimensions, from stance and sentiment detection to topical framing and contradiction analysis. The findings reveal how global media ecosystems construct parallel realities through attribution choices, selective emphasis, and strategic framing, rarely engaging with competing narratives.
Beyond its utility in computational journalism, dnipro provides a foundation for developing more sophisticated methods to detect cross-cultural narrative divergence, identify emergent propaganda strategies, and understand information warfare dynamics. As the global information landscape grows increasingly complex, such resources become essential for technical advancement as well as a broader understanding of narratives that shape our perception of international conflicts. limitations and ethics don’t count towards the page limit
7 Limitations
The Russo-Ukrainian War has not concluded. The dataset, which currently spans from Feb 2022 to August 2024, may omit important narratives and developments resulting from politically significant events thereafter (e.g., foreign policy changes of a powerful geopolitical entity). For the chosen news sources, a wider variety of publications representing NATO nations would be beneficial. Currently, only two of the thirty-two NATO nations are represented (the United Kingdom and the United States of America). There are other media sources like DW in Germany, which are highly referenced globally, so they are an influential factor. Additionally, representation of more countries with ambiguous involvement in the war would add value to a truly global version of the dataset. Potentially interesting countries would be Hungary and Romania, both of which border Ukraine; Belarus, also bordering Ukraine, but known to be an ally of the Russian state; Turkiye, a powerful regional actor, especially in the Black Sea; and Georgia, a NATO contender that continues to maintain reasonably strong relations with Russia.
Another facet limiting dnipro is the imperfect ground-truth annotations for stance, and other discursive phenomena. While this is not a limitation for qualitative analyses or mixed-method approaches, it does hinder nuanced supervised learning directly from the corpus. Our exploratory analyses, while illustrative, are not exhaustive and serve primarily to demonstrate the dataset’s potential rather than to make definitive empirical claims.
8 Ethical Considerations
dnipro contains potentially sensitive content related to an ongoing conflict with intense geopolitical implications. We acknowledge several potential ethical dimensions:
  • Data collection: All content was gathered from publicly available news sources without breaching access restrictions or terms of service. We do not include any copyright-protected content or any private communications.
  • Bias, representation, and inclusion: Our corpus inherently reflects the biases present in global media coverage. The inclusion of opposing narratives—including state-controlled media—is a deliberate choice, to empower comparative analysis. We strongly caution users against interpreting the mere inclusion of a source as endorsement of its content.
  • Potential misuse: While designed for media analysis, computational linguistics, and conflict studies, this resource could also be used for propaganda analysis. And thus, it could also be misused for propaganda analysis. We strongly urge the community to utilize this corpus for constructive research into understanding narrative differences to bridge societal gaps, rather than weaponize any findings.
  • Lost in translation: Machine translation introduces potential distortions, particularly with politically charged terminology and diplomatic language use. We document the details of our translation process. Further, we retain the original articles for enhanced scrutiny and possible resolution by experts in Russian, Ukrainian, and Mandarin Chinese.
  • Contextualization: We provide robust metadata (publication date, source country, outlet ownership) to ensure articles can be interpreted within their proper contexts rather than in isolation.
By openly addressing these considerations, we aim to enable valuable research on contentious discourse while minimizing potential harms. It is our expectation that through the study of such discourse, researchers will be able to bring empirical enhancements to conflict resolution mechanisms. Applications of such datasets are continuously evolving, as is the dataset itself. As such, we earnestly welcome community feedback on additional safeguards.
Appendix A Appendix
A.1 Prompt Design Details
For our quantitative analysis, we used the mistral-small-latest model via the Mistral API to jointly extract stance targets and classify stances, and employed the siebert/sentiment-roberta-large-english model from Hugging Face to estimate sentiment polarity toward named entities. We used GPT-3.5 (via the OpenAI API) to assist with annotation for three tasks: article-level tag extraction, claim identification, and claim type classification. Prompt configurations varied by task. We employed both zero-shot and few-shot strategies, along with system-level instructions and output formatting constraints, following best practices in prompt engineering for language models Saravia (2022).
A.1.1 Annotation for Stance
For stance detection, we used the mistral-small-latest model via the Mistral API. The task was to classify whether a piece of news content expressed a stance FAVOR, AGAINST, or NONE toward a generated target. We followed a combined target generation at first followed by stance classification.  Akash et al. (2025) We ran these experiments on 501 articles that mentioned the Bucha massacre dating from March to April 2022.
The following zero-shot prompt was used to guide the model:
Prompt:
Listing 1: Zero-shot prompt for joint target generation and stance classification.
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a specified target. Analyze the following text, generate the target for this text, and determine its stance toward the generated target. A target should be the topic discussed in the text. The target can be a single word or a phrase, but its maximum length must be 4 words.
If the stance is in favor of the generated target, write FAVOR; if it is against the target, write AGAINST; if it is ambiguous, write NONE. The answer must be exactly one of these three words: FAVOR, AGAINST, or NONE.
Output format:
Target: <target>, Stance: <stance>
A.1.2 Annotation for Sentiment
We used the siebert/sentiment-roberta-large-english model from Hugging Face to perform sentence-level sentiment classification on all news articles. For each entity of interest (e.g., Russia, Ukraine, Bucha), we first extracted all sentences from the article that mentioned the entity, using spaCy’s English language model for sentence segmentation.
Prompt:
What is the sentiment(positive, negative, or neutral) towards {entity} in this context: ’{text}’?
A.1.3 Annotation for Article-level Tags
To annotate articles with high-level thematic tags, we used a zero-shot prompt with GPT-3.5 that instructed the model to extract a set of concise, specific tags summarizing each article’s content. Tags were intended to capture relevant people, locations, military units, political themes, and ideological framings. To ensure consistency and support post-processing, the prompt enforced a structured JSON output format.
Prompt:
You are a geopolitical analyst trained to extract concise topical tags from news articles related to international conflict, especially the Russia–Ukraine war. Your task is to read the content of a single article and extract 5 to 12 relevant tags that capture the article’s key topics, actors, places, events, or narrative themes. Each tag should be 1–4 words long, lowercase unless a proper noun, and specific rather than general. Avoid vague tags like “fake”, “news”, or “war” unless central to the article. Tags may include named individuals, military groups, operations, locations, events, or ideological narratives. Output the result in valid JSON using the following structure:
  "article_123": [
    "bucha incident",
    "sergei lavrov",
    "false flag allegations",
    "media manipulation",
    "un security council"
A.1.4 Annotation for Claim Extraction with Summary and Claim Types
We used a structured few-shot prompt with GPT-3.5 to extract and classify claims from news articles. Each prompt instructed the model to identify distinct claims within an article, summarize them neutrally, and assign a claim_type based on a predefined taxonomy. Claims could span one or more sentences and were returned in a structured JSON format to support post-processing.
Prompt:
You are a geopolitical analyst trained to extract claims from news articles about the Russia–Ukraine war. Your task is to read the content of a single article and extract each distinct claim it makes. For each claim, return the original sentence(s), a short neutral summary, and a claim_type. Claims may span multiple sentences. Use only the article content field for analysis. The claim_type must be one of the following: Factual, Attributed, Disputed, Reported, Command/Policy, or Analytical/Assessment. Output the result in valid JSON using the following structure:
  "article_123": {
    "claim_1": {
      "sentence":
        "Ukrainian Side Reports Multiple
        Casualties in Kyiv Airstrike.",
      "summary":
        "Ukraine reports casualties in
        Kyiv airstrike.",
      "claim_type": "Attributed"
    "claim_2": {
      "sentence":
        "Zelensky Orders: ’Ensure NATOs
        Long-Term Presence in the Black
        Sea’.",
      "summary":
        "Zelensky orders long-term NATO
        presence in the Black Sea.",
      "claim_type": "Command/Policy"
A.2 Use Case Experiment Details
Figures 4(a) and 4(b) illustrate 5-day interval favor stance percentages toward Russia and Ukraine across news source countries during March–April 2022.
(a) favor stance towards Russia (by source country).
(b) favor stance towards Ukraine (by source country).
Figure 4: 5-day interval favor stance percentages toward Russia and Ukraine (during the Bucha massacre, March–April 2022). News outlets are grouped by their national affiliations.

### v2

@@
 Abstract
-We introduce dnipro, a novel longitudinal corpus of 246K news articles documenting the Russo-Ukrainian war from Feb 2022 to Aug 2024, spanning eleven media outlets across five nation states (Russia, Ukraine, U.S., U.K., and China) and three languages (English, Russian, and Mandarin Chinese).
-This multilingual resource features consistent and comprehensive metadata, and multiple types of annotation with rigorous human evaluations for downstream tasks relevant to systematic transnational analyses of contentious wartime discourse.
-dnipro’s distinctive value lies in its inclusion of competing geopolitical perspectives, making it uniquely suited for studying narrative divergence, media framing, and information warfare.
-To demonstrate its utility, we include use case experiments using stance detection, sentiment analysis, topical framing, and contradiction analysis of major conflict events within the larger war.
-Our explorations reveal how outlets construct competing realities, with coverage exhibiting polarized interpretations that reflect geopolitical interests.
-Beyond supporting computational journalism research, dnipro provides a foundational resource for understanding how conflicting narratives emerge and evolve across global information ecosystems.
+We present dnipro, a corpus of 246K news articles from the Russo-Ukrainian war (Feb 2022 – Aug 2024) spanning eleven outlets across five nation-states (Russia, Ukraine, U.S., U.K., China) and three languages.
+The corpus features comprehensive metadata and human-evaluated annotations for stance, sentiment, and topical framing, enabling systematic analysis of competing geopolitical narratives.
+It is uniquely suited for empirical studies of narrative divergence, media framing, and information warfare.
+Our exploratory analyses reveal how media outlets construct incompatible realities through divergent attribution and topical selection without direct refutation of opposing narratives.
+dnipro empowers empirical research on narrative evolution, cross-lingual information flow, and computational detection of implicit contradictions in fragmented information ecosystems.
+Keywords: multilingual corpus, media framing, geopolitical narratives, longitudinal analysis, information warfare
+[…]
 A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War
-Dikshya Mohanty, Taisiia Sabadyn, Jelwin Rodrigues, Chenlu Wang, Abhishek Kalugade, Ritwik Banerjee Stony Brook University Correspondence: {dimohanty, jelrodrigues, rbanerjee}@cs.stonybrook.edu
-1 Introduction
-The Russo-Ukrainian war, which escalated on February 24, 2022 with the declaration of a "special military operation" by the Russian government, [^1] Within 48 hours, media in Russia were banned from describing the developments as “war” AFP (2022).
+Dikshya Mohanty1, Taisiia Sabadyn1, Jelwin Rodrigues1,
+Chenlu Wang1, Abhishek Kalugade1, Ritwik Banerjee1,2
+1Department of Computer Science, 2AI Innovation Institute
+Stony Brook University, New York
+{dimohanty, jelrodrigues, chenlwang, rbanerjee}@cs.stonybrook.edu
+{taisiia.sabadyn, abhishek.kalugade}@stonybrook.edu
+Abstract content
+1.  Introduction
+The Russo-Ukrainian war, which escalated on February 24, 2022 with the declaration of a “special military operation” by the Russian government, [^1] Within 48 hours, media in Russia were banned from describing the developments as “war” AFP (2022).
 has significantly impacted the global geopolitical landscape Derviş (2023); Silvertsen (2024).
@@
 Yet, despite the centrality of media and narrative framing, there remains a lack of comprehensive datasets for longitudinal analysis of how nations have framed the war for their domestic and international audiences.
-We introduce a collection of Diverse Narratives and International Perspectives on the Russo-Ukrainian Offensive (dnipro), a comprehensive longitudinal corpus of news coverage from February 2022 to August 2024222An expanded version is being developed, which will include news coverage up to February 2025 and provide a longitudinal 3-year corpus of this protracted war.—the longest time span of any multinational or multilingual news corpus on the Russia-Ukraine war.
+We introduce a collection of Diverse Narratives and International Perspectives on the Russo-Ukrainian Offensive (dnipro), a comprehensive longitudinal corpus of news coverage from February 1, 2022 to August 31, 2024 — the longest time span of any multinational or multilingual news corpus on the Russia-Ukraine war.
+The code-base and supporting materials are publicly available.
 The corpus encompasses eleven influential media outlets and news agencies aligned with five primary geopolitical actors with strategic interests in the conflict.
@@
 (2016) and Roman et al.
-(2017), the sources include media from direct participants—Russia and Ukraine—and influential global powers with economic, diplomatic, or military involvement: USA, UK, and China.
-The inclusion of state-affiliated media from Russia and Ukraine capture opposing conflict narratives, while those from China, the USA, and the UK represent the narratives projected by nations with significant stakes in the conflict’s outcome, and significant ability to influence that outcome Zhang et al.
+(2017), the sources include media from direct participants—Russia and Ukraine—and influential global powers with economic, diplomatic, or military involvement: the USA, the UK, and China.
+The inclusion of state-affiliated media from Russia and Ukraine captures opposing conflict narratives that play a pivotal role in shaping assessments under uncertain information Saletta et al.
+(2020).
+Media from China, the USA, and the UK represent perspectives of nations with substantial geopolitical and economic interests in the conflict’s trajectory, as well as the military, diplomatic, and economic leverage to influence its outcome Zhang et al.
 (2016); Thussu et al.
 (2017).
-dnipro offers rich annotations for entity-level stance, sentiment, and topical framing, enabling event tracking and other longitudinal analyses across geopolitical perspectives.
-The dataset does not attempt to verify the accuracy of any claims; rather, it provides a resource for studying how wartime narratives are constructed and disseminated by major geopolitical actors.
+These diverse perspectives and sources enable more comprehensive analysis of complex developments by helping analysts overcome the limitations of single-narrative framings Werd (2018).
+dnipro offers rich annotations for topical framing, entity-level stance, and entity-level sentiment, enabling event tracking and other longitudinal analyses across geopolitical perspectives.
+The dataset does not attempt to verify the accuracy of any claims; rather, it provides a resource to study how wartime narratives are constructed and disseminated by major geopolitical actors.
 [^2] This aligns with established methodologies in discourse analysis Entman (1993); Fairclough (2010) examining how language and framing construct social reality.
-Further, Gamson and Modigliani (1989) showed how this view offers a powerful tool to understand the competition that shapes public understanding of contentious events.
+Gamson and Modigliani (1989) demonstrated how this approach reveals the competition shaping public understanding of contentious events.
 The corpus is uniquely suited to such studies, as it comprises 246,229 articles across 11 media sources and 3 languages, spanning 5 diverse nations, over a period of 31 months.
@@
   • (1)
-    a longitudinal multinational and multilingual corpus of news coverage spanning the first 30 months of the Russia-Ukraine war;
+    a longitudinal, multinational, and multilingual corpus of news coverage spanning the first 31 months of the Russia-Ukraine war;
   • (2)
-    annotations supporting comparative analyses of named entities, stance, and framing; and
+    inclusion of diverse media, enabling cross-cultural examination of conflict narratives; and
   • (3)
-    the inclusion of diverse media sources enabling cross-cultural examination of geopolitical narratives.
-Next, we place dnipro within the landscape of existing conflict-related corpora and highlight its distinct contributions to the field (§ 2), and describe how our design and construction choices (§ 3) make it a powerful resource for large-scale analyses of wartime narratives.
-We employ dnipro in several applications, using state-of-the-art language understanding systems as well as additional annotations (§ 4 and § 5), demonstrating its utility.
-2 Related Work
+    annotations for comparative analyses of named entities, sentiment, stance, and topical framing.
+Next, we place dnipro within the landscape of existing conflict-related corpora to highlight its distinct contributions to the field (§ 2), and describe how our design and construction choices (§ 3) make it a powerful resource for large-scale analyses of wartime narratives.
+We employ dnipro in several applications, using state-of-the-art natural language understanding systems and additional annotations (§ 4 and § 5), demonstrating its utility.
+2.  Related Work
 Research on wartime media coverage relies on datasets capturing narratives across national and institutional contexts, where competing claims and propaganda techniques pose unique challenges.
@@
 As a reflection of diverse goals, existing resources vary in scope, modality, and annotations.
-Twitter & Reddit-based datasets are among the largest.
+Datasets based on Twitter/X and Reddit are among the largest.
 Chen and Ferrara (2023) aggregate over 570 million tweets in English, Russian, and Ukrainian to study information propagation and influence operations, while Zhu et al.
@@
 The VoynaSlov collection Park et al.
-(2022) bears similarities to our work through sentence-level media framing labels, though they focus solely on Russian media.
+(2022) bears similarities to our work through sentence-level media framing labels, though it focuses solely on Russian media.
 Alyukov et al.
-(2023) introduce one of the first cross-platform social media datasets, integrating 1.7 million posts to enable propaganda and public opinion analyses.
+(2023) introduce one of the first cross-platform social media datasets, integrating 1.7 million posts to enable the analysis of propaganda and public opinion.
 Telegram-specific corpora include 4.1 million posts with pro/anti-Kremlin channel tags Bawa et al.
@@
 (2023).
-Hakimov and Cheema (2024) introduce a corpus of 1.5 million tweets in 60 languages with automated sentiment, stance, and named entity annotations, offering the widest multilingual social media dataset.
+Hakimov and Cheema (2024) introduce a corpus of 1.5 million tweets in 60 languages with automated annotations for sentiment, stance, and named entities.
+Their corpus is one of the widest multilingual social media datasets currently available.
+News-based corpora are comparatively sparse.
 The RUWA corpus Khairova et al.
@@
 dnipro supersedes it by orders of magnitude in coverage, topics, multilinguality, temporality, and size.
-Ibrahim et al.
-(2025) offer longitudinal media analysis of two years of news data, but from nations with auxiliary involvement.
-Annotation practices differ widely: manual schemata for framing Park et al.
+A longitudinal media analysis corpus is provided by Ibrahim et al.
+(2025), with two years of news data, but its coverage is restricted to nations with auxiliary involvement in the Russo-Ukrainian war.
+Annotation practices across existing resources differ widely: manual schemata for framing Park et al.
 (2022), channel stance Bawa et al.
 (2025), and human rights violations Nemkova et al.
-(2023); automated annotations for large-scale labeling Susanne Pohl et al.
+(2023); as well as automated large-scale labeling Susanne Pohl et al.
 (2023); Hakimov and Cheema (2024).
 Most resources rely on social media identifiers that degrade quickly, impeding reproduction.
-Few provide longitudinally contiguous, multilingual, news-based content annotated for multiple tasks.
-dnipro extends existing corpora by focusing on transnational and multilingual news coverage – underrepresented in prior social media-centered work – enabling comparative analyses of state-aligned narratives across geopolitical actors.
-It reflects news coverage as disseminated, without alteration, serving as a neutral resource for empirical studies.
-Our design choices spanning source selection, temporal coverage, and structured annotations make dnipro a powerful resource for large-scale automated analysis of wartime narratives (§ 5).
-3 The dnipro Corpus
-Our dataset is constructed through a systematic multi-stage pipeline targeting news outlets and state-affiliated agencies selected on the basis of three criteria: (1) global influence and prominence within national media ecosystems, (2) sustained coverage of the Russo-Ukrainian war, and (3) discernible editorial positioning ranging from independent journalism to state-aligned messaging.
+dnipro addresses these gaps by providing longitudinally contiguous, multilingual news coverage from five primary geopolitical actors (Russia, Ukraine, U.S., U.K., and China), with comprehensive annotations for multiple downstream tasks.
+Unlike social media-focused datasets optimized for volume and engagement patterns, dnipro enables systematic comparative analyses of state-aligned narrative framing across languages, nations, and outlets, over a duration of 2 years and 7 months, making it a unique and powerful resource for empirical studies on the creation and dissemination of competing wartime realities (§ 5).
+3.  The dnipro Corpus
+We present dnipro, a longitudinal corpus of 246,229 news articles spanning 31 months of the Russo-Ukrainian war (February 1, 2022 – August 31, 2024).
+Our corpus captures competing narratives from 11 outlets across five nation states (Russia, Ukraine, U.S., U.K., and China) and three languages (English, Russian, and Mandarin Chinese).
+Its key descriptive statistics appear in Table 1.
+Each article includes structured metadata following a unified schema (shown in Table 2) designed to facilitate large-scale longitudinal and comparative analyses across sources, languages, and geopolitical contexts.
+dnipro is constructed through a systematic multi-stage pipeline targeting news outlets and state-affiliated agencies selected on the basis of three criteria: (i) global influence and prominence within national media ecosystems, (ii) sustained coverage of the Russo-Ukrainian war, and (iii) discernible editorial positioning ranging from independent journalism to state-aligned messaging.
 To the extent possible, access to full article archives was obtained through public RSS feeds, institutional subscriptions, or source-specific data collection pipelines.
 However, not all media outlets provided consistent or unrestricted access to their archives, and some selections reflect practical constraints in addition to topical relevance.
-3.1 Data Acquisition
-dnipro is a corpus spanning 31 months of daily data collection, from February 1, 2022 to August 31, 2024.
+We first detail our collection methodology (§3.1), then describe the structure and access information for the data records (§3.2).
+Corpus Scale
+Total articles 246,229
+Avg. no. of articles/day 261.1
+Title length (avg.) 11.7 words
+Article length (avg.) 332.5 words
+Coverage
+Temporal span Feb 1, 2022 – Aug 31, 2024
+No. of news sources 11 sources
+Geopolitical entities Russia, Ukraine, USA, UK, China
+Linguistic Composition
+Languages covered English, Russian, Mandarin Chinese
+Non-English articles 61,637 (25.03%)
+After translation (of non-English sources to English).
+Table 1: dnipro at a glance: scale, temporal and geographic coverage, and multilingual composition.
+3.1.  Data Acquisition
+dnipro comprises 31 months of continuous daily data collection (Feb.
+1, 2022 – Aug.
+31, 2024).
 To identify articles with substantive coverage of the Russo-Ukrainian conflict, we employ a tiered strategy.
@@
 To ensure quality control, we manually validate stratified random samples across sources, languages, and time periods throughout the acquisition process.
-Total no. of articles 246,229
-Title length (avg.) 11.7 words
-Article length (avg.) 332.5 words
-No. articles/day (avg.) 261.1
-No. of sources 11
-Languages covered N/A
-Temporal span Feb 1, 2022 - Aug 31, 2024
-No. of non-English articles 61,637 (25.03%)
-Geopolitical entities Russia, Ukraine, China, USA, UK
-After translation (of non-English sources to English).
-Table 1: Key descriptive statistics of dnipro.
-English-language “Western” media from the US and the United Kingdom, namely The Financial Times (FT) and Cable News Network (CNN), were included as part of the Western bloc, reflecting the perspectives of countries with significant political, economic, and military involvement in the conflict.
-These outlets serve domestic and international audiences and are central to understanding how Western-aligned narratives are constructed and disseminated.
+Field Description
+Identifiers & Source
+article_id Unique UUID v4 identifier
+url URL of the article
+source_name Normalized source name (e.g., Xinhua)
+source_country Country of origin (ISO 3166-1 alpha-3)
+Content & Language
+title Article headline
+article_length Word count (post-processing)
+language Article language (ISO 639-1)
+Temporal Information
+publication_date Publication date (ISO 8601)
+collection_date Date of URL access (ISO 8601)
+Table 2: Core metadata schema for dnipro articles.
+English-language “Western” media from the United States and the United Kingdom, namely The Financial Times (FT) and Cable News Network (CNN), were included as part of the Western bloc, reflecting the perspectives of countries with significant political, economic, and military involvement in the conflict.
+These outlets serve domestic and international audiences, and offer central platforms to help us understand how Western-aligned narratives are constructed and disseminated.
 No additional filtering of articles is done by search queries as these media houses offer dedicated sections for content related to the Russo-Ukrainian war.
@@
 […]
-The Ukrainian perspective is represented by articles from Censor.net, The Kyiv Independent (KI), and European Pravda (EP), all of which provide English-language versions.
-While KI and EP shape Ukraine’s global image, Censor influences the domestic audience.
-Their reporting styles vary—from Censor’s rapid updates to EP’s in-depth opinions and interviews.
-As sources targeting the global Anglosphere, KI and EP offer sections dedicated to the Russo-Ukrainian War: “War” and “War with Russia”, respectively.
-Censor’s news reporting approach largely covers the war, evidenced by the majority fraction of Ukrainian war-related tags in their “News” section.
+The Ukrainian perspective is represented by articles from Censor.net, The Kyiv Independent, and European Pravda, all of which provide English-language versions.
+While Censor.net influences Ukraine’s domestic audience, the other two shape her global image.
+Their reporting styles vary—from Censor’s rapid updates to the in-depth opinions and interviews found in European Pravda.
+As sources targeting the global Anglosphere, The Kyiv Independent and European Pravda offer sections dedicated to the Russo-Ukrainian War, labeled respectively with “War” and “War with Russia”.
+The news reporting approach of Censor.net largely covers the war, evidenced by the majority fraction of Ukrainian war-related tags in their “News” section.
 As such, no keyword search queries are applied to any of the Ukrainian sources.
-The collection from these courses amount to 19,706 articles from Kyiv Independent, 2,212 from European Pravda, and 31,748 from Censor.
+The collection from these sources amounts to 19,706 articles from The Kyiv Independent, 2,212 from European Pravda, and 31,748 from Censor.net.
 The Russian media sources in our dataset comprise English articles from TASS and Sputnik, and Russian-language articles from Izvestia.
@@
 We collect 60,002 and 53,139 articles from TASS and Sputnik respectively, and 50,485 articles from Izvestia.
-article_id A unique UUID v4 identifier for the record.
-url The persistent URL from which the article was sourced.
-publication_date The article’s publication date in ISO 8601 format (yyyy-mm-dd).
-source_name The normalized name of the news outlet (e.g., Xinhua).
-source_country The country of origin for the news outlet, in ISO 3166-1 alpha-3.
-language The article language, in ISO 639-1.
-title The article’s title (i.e., headline).
-article_length The number of words in the cleaned article text (post processing).
-collection_date The date (in ISO 8601 format) when the URL was accessed.
-Table 2: Core metadata fields for each article.
-3.2 Metadata Schema and Statistics
-The core of dnipro is a collection of rich, consistent metadata records for 246,229 news articles.
-Its key descriptive statistics are shown in Table 1.
-Each record is described by a unified schema (described in Table 2) designed to facilitate large-scale longitudinal and comparative analyses.
-The overall composition of the natural language data—both in terms of the number of news articles and their sizes—across individual news sources and aggregated by geopolitical entities, is presented in Figure 1.
-3.3 Data Records
-dnipro is publicly available on Zenodo under a CC-BY-NC 4.0 license.
-[^4] The data, along with all processing code and annotation prompts, will be released upon acceptance.
-It is structured for ease of use and distributed as a single, compressed tar.bz2 archive containing a primary metadata file, annotation files, and a documentation file.
-[^5] We do not directly include the texts due to copyright restrictions.
-The provided URLs serve as persistent identifiers enabling reproduction of dnipro, like most social media datasets.
-If an article becomes unavailable, the collection_date and url enable retrieval via the Internet Archive’s WayBack Machine.
+Figure 1 shows the distribution of articles and their average lengths across individual news sources and aggregated by geopolitical entities, providing an overview of the varying scales and styles of coverage across geopolitical perspectives.
+3.2.  Data Records
+The dnipro corpus is available at no cost for non-commercial use under the CC-BY-NC 4.0 license, [^3] The complete corpus is available on Zenodo: doi.org/10.5281/zenodo.18470677 and the corresponding processing scripts and annotation prompts are available on GitHub.444github.com/dikshyam/dnipro-codebase It is structured for ease of use and distributed as separate downloadable files, comprising a primary metadata file, annotation files, and a documentation file.
+Due to copyright restrictions, we cannot directly include the texts.
+The provided URLs, however, serve as persistent identifiers enabling reproduction of dnipro, like most social media datasets.
+If an article becomes unavailable, the collection_date and url fields enable retrieval via the Internet Archive’s WayBack Machine.555web.archive.org
   1. 1.
-    metadata.parquet contains the complete set of 246,229 metadata records as described in § 3.2.
+    metadata.parquet contains the complete set of 246,229 metadata records as described in Table 2.
 The Parquet format was chosen for its efficient compression and columnar storage, which enables fast querying and analysis using popular data science tools.
@@
   3. 3.
+    topical-frames.parquet provides one row per article_id with its corresponding topical frames anchored in categories proposed by Habermas (1991) (Table 3).
+  4. 4.
     sentiments.parquet provides sentiment scores for select entities and events across multiple news sources.
-For politically important and frequently occurring entities (e.g., Ukraine, Russia, Zelensky, Kherson), these scores are aggregated monthly.
-For entities tied to key events within this war (e.g., Bucha, Kursk, Zaporizhzhia), they are aggregated daily.
-  4. 4.
-    topical-frames.parquet provides one row per article_id with its corresponding topical frames anchored in the nine-category schema described in Table 3.
+For politically important and frequently occurring entities (e.g., “Ukraine”, “Russia”, “Zelensky”, “Kherson”), these scores are aggregated monthly.
+For entities tied to key events within this war (e.g., “Bucha” or “Zaporizhzhia”), they are aggregated daily.
+The events used in this work are described in § 5 (Table 4).
   5. 5.
@@
 We encourage task-appropriate partitions instead.
-4 Data Annotations
+4.  Data Annotations
 We now describe the enrichment of raw article metadata with annotations.
@@
 The process involves:
-  1. 1.
   • metadata and text normalization;
-  2. 2.
   • translation of non-English text to English;
-  3. 3.
   • labeling named entities;
-  4. 4.
   • sentiment analysis;
-  5. 5.
   • stance analysis;
-  6. 6.
   • topical framing; and
-  7. 7.
   • rigorous quality control with human evaluation.
-4.1 Text Processing for Annotation
+4.1.  Text Processing for Annotation
 To prepare the corpus for annotation and analyses, we apply standardized processing steps to all data, across sources and languages.
 This ensures consistency across media sources with varying conventions in structure, formatting, and metadata.
-Publication dates are parsed and normalized to ISO 8601 standard (yyyy-mm-dd), and multiple metadata fields—e.g., “publication date”, “published on”—are mapped to metadata fields in our unified schema (see Table 2).
-Text content is stripped of non-informative boilerplate elements like location or date prefixes commonly found in news articles (e.g., “KYIV, March 3 –”), inline footnotes, bylines, and repeated source disclaimers.
+Publication dates are parsed and normalized to ISO 8601 standard (yyyy-mm-dd), and multiple metadata fields—e.g., “publication date”, “published on”—are mapped to metadata fields in our unified schema (as shown in Table 2).
+The article text content is stripped of boilerplate elements like location or date prefixes commonly found in news articles (e.g., “KYIV, March 3 –”), inline footnotes, bylines, and repeated source disclaimers.
 Duplicate articles are removed, to avoid redundancy in print syndication and multilingual reprints.
@@
 Frame Definition
-economy Macroeconomic issues and markets: sanctions/trade, inflation, growth, jobs, taxes, energy/commodity prices.
-identity Group identity and belonging: nationality, ethnicity, religion, language, or gender; discrimination or identity politics.
-morality Explicit ethical/value framing: moral duty, fairness, dignity, humanitarian arguments.
-legal Courts and legal processes: rulings, lawsuits, prosecutions, designations, warrants, legal rights or entitlements.
-policy Concrete government outlook or outreach, rules, bills, decrees, program design and details (what a policy does).
-politics elections, party/elite competition, horse-race strategy, reshuffles, coalition dynamics.
-public opinion Polls, surveys, attitudes and protests as expressions of public sentiment.
-security War/defense, violence/crime, policing, border enforcement, terrorism, cyberattacks.
-welfare Benefits and human services: healthcare, pensions, housing aid, unemployment, assistance to refugees or internally displaced people.
-Table 3: Topical frame definitions in dnipro: a significant abridgment of the Refugees and Migration Framing Vocabulary Yu and Fliethmann (2022).
-4.2 Machine Translation
+economy 1 Macroeconomic issues and markets: sanctions/trade, inflation, growth, jobs, taxes, energy/commodity prices.
+identity 3 Group identity and belonging: nationality, ethnicity, religion, language, or gender; discrimination or identity politics.
+morality 2 Explicit ethical/value framing: moral duty, fairness, dignity, humanitarian arguments.
+legal 1 Courts and legal processes: rulings, lawsuits, prosecutions, designations, warrants, legal rights or entitlements.
+policy 1 Concrete government outlook or outreach, rules, bills, decrees, program design and details (what a policy does).
+politics 1 Political proceedings and party competition: elections, horse-race strategy, reshuffles, coalition dynamics.
+public opinion 1 Public attitudes and moods: polls, surveys, and protests as expressions of public sentiment and interest.
+security 1 War/defense, violence/crime, policing, border enforcement, terrorism, cyberattacks.
+welfare 1 Benefits and human services: healthcare, pensions, housing aid, unemployment, assistance to refugees or displaced people.
+Table 3: Topical frames in dnipro: an abridgment of the Refugees and Migration Framing schema Yu and Fliethmann (2022), with its utilitarian 1, moral-universal 2, and identity-related 3 categories derived from the discourse ethics of Habermas (1991).
+4.2.  Machine Translation
 A quarter of dnipro consists of Mandarin Chinese and Russian texts, (zh and ru, respectively, in ISO 639-1).
@@
 (2021), and NLLB-200-Distilled Team (2022).
-Since there were no readily available explicit benchmarks for zh-to-en or ru-to-en translations, we select a random set of ten articles from Xinhua, evaluated by two native speakers of Mandarin Chinese for semantic fidelity and fluency.
+Since there were no readily available explicit benchmarks for zh-to-en translation of war or conflict narratives, we select a random set of ten articles from Xinhua, evaluated by two native speakers of Mandarin Chinese for semantic fidelity and fluency.
 The evaluators worked independently to score the output of these models on 1-5 Likert scale for fidelity and fluency, and M2M100-418M received the highest scores from both evaluators on both criteria.
 So, we employ this model for all zh-to-en translations.
-Russian articles were similarly translated using the ru-to-en path of M2M100-418M.
-4.3 Annotation Methodology
-To enable fine-grained analysis of key actors and locations, we perform Named Entity Recognition (NER) across the entire corpus (after translation to English).
-This annotation is carried out using the spaCy library Honnibal et al.
-(2020) and its pretrained model en_core_web_sm, identifying entities and their types according to the OntoNotes 5.0 taxonomy Weischedel et al.
+Russian articles from Izvestia are similarly evaluated by native Russian speakers, and translated using the ru-to-en path of M2M100-418M.
+4.3.  Annotation Methodology
+Named Entity Recognition: To enable fine-grained analysis of key actors and locations, we perform named entity recognition (NER) across the entire corpus (after translation to English).
+This annotation is carried out using spaCy Honnibal et al.
+(2020) and its pretrained model en_core_web_sm, identifying entities and their types according to the OntoNotes 5.0 taxonomy (ontonotes).
+This identifies 85,155 distinct named entity mentions, and provides the foundational layer for downstream tasks such as tracking the prominence of specific actors over time, temporal studies of media focus across entities and events, and entity-specific sentiment and stance analyses.
+Sentiment Analysis: To capture the sentiment surrounding the most prominent named entities, we provide sentence-level sentiment annotation using a pretrained model: sentiment-roberta-large-english Hartmann et al.
 (2023).
-This process results in the identification of 85,155 distinct named entity mentions, and provides a foundational layer for tasks such as tracking the prominence of specific actors over time, temporal studies of media focus across entities and events, and entity-specific sentiment and stance analyses.
+For each article, sentences containing such entities (e.g., Kursk, Bucha, Kherson, Zelensky, Ukraine, Russia) are identified using spaCy, and passed to this classifier.
+The model outputs are then mapped to continuous sentiment scores, which are then aggregated on a daily and monthly basis at both the source and country level using source–country relations.
+These annotations thus provide time series data revealing how sentiment toward specific actors and events evolves across geopolitical perspectives.
 Event (Date range) # articles
@@
 Table 4: The four conflict events analyzed in our use case experiments, selected based on global coverage and the potential to generate narrative divergence.
-To capture the sentiment of some of the most prominent named entities, we provide sentence-level sentiment annotation using a pretrained model: sentiment-roberta-large-english Hartmann et al.
-(2023).
-For each article, sentences containing such entities (e.g., Kursk, Bucha, Kherson, Zelensky, Ukraine, Russia) are identified using spaCy, and passed to this classifier.
-The model outputs are then mapped to continuous sentiment scores, which are then aggregated on a daily and monthly basis at both the source and country level using source–country relations.
-These annotations thus provide time series data revealing sentiment trends for each entity.
-The dataset includes stance annotations obtained by few-shot instruction prompting with Phi-4 14B model  Microsoft (2024), identifying the stance towards a specified target.
+Stance Detection: The dataset includes stance annotations obtained by few-shot instruction prompting with Phi-4 14B model  Microsoft (2024), identifying the stance towards a specified target.
 The input comprises a sentence and the immediate context in its news article.
-As part of the input, we provide a task instruction with a few examples, along with (i) a passage for context resolution, (ii) a focal sentence, and (iii) the target.
+To Phi-4, we provide the task instruction with a few examples, along with (a) a passage for contextual information, (b) a focal sentence, and (c) the target.
 The model is prompted to output a compact, formatted response:
@@
     Rationale: <a short justification grounded in the given text>
-For topical framing, we again use few-shot prompting with Phi-4 14B model, where the input is an instruction describing the task, along with label definitions and 1-2 examples per label.
-Then, for a given instance – an article’s title, subtitle (if present) and content – we prompt the model to generate a json object containing the list of applicable topical frames, and associated confidence scores.
-4.4 Human Evaluation
-To evaluate the reliability of our automated annotations, we conduct human evaluations of the stance and topical frame annotations.
-Frame PABAK Confidence
-Human/Human Human/LLM Human LLM
-Economy 0.70 (84.8%) 0.736 0.84 0.93
-Identity 0.74 (86.8%) 0.783 0.79 0.92
-Legal 0.85 (92.4%) 0.783 0.74 0.91
-Morality 0.75 (87.3%) 0.750 0.79 0.91
-Policy 0.55 (77.7%) 0.526 0.78 0.89
-Politics 0.48 (74.1%) 0.425 0.79 0.92
-Public Opinion 0.87 (93.4%) 0.865 0.80 0.89
-Security 0.69 (84.3%) 0.587 0.85 0.93
-Welfare 0.81 (90.4%) 0.783 0.80 0.87
-Average 0.71 (85.7%) 0.693 0.798 0.909
-Table 5: Inter-annotator agreement and confidence by topical frame.
-Pabak columns report human inter-annotator agreement (percent agreement in parentheses) and human–model agreement.
-Confidence columns report average confidence scores from human annotators and LLM (Phi-4).
-Topical Frames:
-Two expert annotators independently label a stratified random sample of 200 articles, providing nine binary labels with corresponding [0,1] confidence scores for each article.
-Significant class imbalance is a common challenge in multi-label annotation, and the inter-annotator agreement must consider the high base rate agreement where annotators agree on a label being absent from an article.
-The average agreement on label absence, over all nine topical frames, accounts for 65.82% of our evaluation sample, with the positive class (for any label) representing a minority of instances.
-Due to this prevalence bias, traditional measures like Cohen’s or Fleiss’ kappa are overly pessimistic reliability measures for our data.
-For a more accurate assessment that accounts for this imbalance, we report the prevalence-adjusted bias-adjusted kappapabak Byrt et al.
-(1993) – which is especially crucial for evaluation of consensus on the rarer labels.
-The inter-annotator agreement results, measured by pabak, are presented in Table 5 and demonstrate substantial agreement on average (0.71, with 85.7% agreement).
-Three topical frame categories — Welfare, Public Opinion, and Legal — reveal near-perfect agreement (>> 0.8).
-Overall, seven of the nine labels achieve pabak scores $\geq$ 0.7, indicating a consistent and shared understanding of the annotation framework.
-The two remaining classes, Policy and Politics, have moderate agreement, as they represent more abstract and often overlapping conceptual domains.
-We also examine the correlation between the average confidence scores reported by the human experts and those reported by Phi-4.
-The framing categories are complex and sometimes overlapping, resulting in a weak positive correlation $\rho$ = 0.37.
-Implications for dataset use.
-The strong agreement validates the overall reliability of our annotation schema.
-However, we recommend that users consider the tiered nature of this reliability when designing their experiments:
-  • The seven frames with pabak > 0.69 are suitable for fine-grained analysis, including temporal tracking and cross-source comparison.
-  • The Policy and Politics labels yield lower agreement due to their inherent conceptual overlap.
-We recommend these labels be used for high-level topic modeling or as supplementary features in larger models, with the understanding that they represent broader and possibly intersecting, categories of discourse.
-  • The human-annotated sample of 200 documents, while modest in size, provides a high-quality validated benchmark.
-It was strategically selected through stratified sampling to ensure longitudinal representation across sources and languages, making it suitable for evaluating the performance of topical frame classification systems on real-world data on contentious wartime narratives.
+Topical Framing: For topical framing, we use the same few-shot prompting with Phi-4 14B (Microsoft (2024)), where the input is an instruction describing the task, along with topical frame label definitions (shown in Table 3) and 1-2 examples per label.
+Then, for a given instance—comprising an article’s title, subtitle (if present) and content—we prompt the model to generate a json object containing the list of applicable topical frames and their corresponding confidence scores.
+This supports analysis of how different sources emphasize humanitarian, military, diplomatic, or economic dimensions of the conflict.
+$4.$  Human Evaluation
+To assess annotation quality, we conduct rigorous human evaluation on both stance detection and topical framing.
+Results demonstrate that automated labeling substantially agrees with expert judgment (pabak = 0.71 for topical framing, Fleiss’ $\kappa$ = 0.643 for stance detection), validating their reliability for coarse-grain analyses while identifying specific categories that warrant careful interpretation.
 Annotator-pair Agreement (%) Cohen’s $\kappa$
@@
 $\bullet\;\;$ Mean pairwise Cohen’s $\kappa$: 0.656 (range: 0.58-0.77)
-Table 6: Inter-annotator agreements for stance labels, measured by percentage of agreement, pairwise Cohen’s $\kappa$, and Fleiss’ $\kappa$.
-Stance Annotations:
+Table 5: Annotation agreement on stance labels: agreement percentage, pairwise Cohen’s $\kappa$, and Fleiss’ $\kappa$.
+4.4.1.  Stance Annotations
 The stance annotation task identifies whether a sentence (with context) expressed a stance (favor or against) towards any entity, or if no stance was present (none).
@@
 We evaluate their inter-annotator agreement, and then also with phi-4 as the LLM annotator.
-Table 6 presents the pairwise and overall inter-annotator agreements, which indicate strong human inter-annotator agreement ($\kappa$ = 0.77).
-The LLM annotator achieved moderate agreement with both ($\kappa$ = 0.62 and 0.58), showing consistent but less-than-human reliability.
-Fleiss’ $\kappa$ = 0.643, however, suggests that the stance annotations are sufficiently reliable for large-scale analysis, despite the gap between human and automated annotation quality for this task.
-Figure 2: Topical framing by source, with % share of articles for each label (confidence >> 0.5).
-Source abbr.: Financial Times (FT), Kyiv Post (KP), Sputnik (SP), European Pravda (EP), Global Times (GT), China Daily (CD), Xinhua (XH).
-Labels: (Sec)urity, (Pol)itics, (Econ)omy, (Welf)are, (Ident)ity, (Moral)ity, and (Pub)lic (Op)inion.
-5 Use Case Experiments
-To demonstrate the analytical potential of dnipro, we present several studies that leverage its unique multi-perspective structure.
-These experiments are intended as use case demonstrations of the dataset’s capacity to support diverse research questions in computational journalism, political science, and discourse analysis.
+Results in Table 5 present the pairwise and overall inter-annotator agreements, demonstrating strong human inter-annotator agreement ($\kappa$ = 0.77) and moderate LLM-human agreements ($\kappa$ = 0.62 and 0.58), showing consistent but less-than-human reliability.
+The overall Fleiss’ $\kappa$ = 0.643 indicates that the stance annotations are substantially reliable for large-scale analysis.
+While the LLM shows a consistent gap with human judgment, this is expected for nuanced pragmatic language tasks like stance detection.
+The annotations remain valuable for trend analysis and comparative studies across sources—such as identifying systematic differences in how Russian vs.
+Ukrainian sources frame political leadership or specific events—while nuanced high-stakes stance analyses may benefit from human review.
+Frame pabak Confidence
+Human All Human Phi-4
+High reliability (pabak > 0.8)
+Public Opinion 0.87 (93.4%) 0.865 0.80 0.89
+Legal 0.85 (92.4%) 0.783 0.74 0.91
+Welfare 0.81 (90.4%) 0.783 0.80 0.87
+Substantial reliability (pabak 0.69 - 0.8)
+Morality 0.75 (87.3%) 0.750 0.79 0.91
+Identity 0.74 (86.8%) 0.783 0.79 0.92
+Economy 0.70 (84.8%) 0.736 0.84 0.93
+Security 0.69 (84.3%) 0.587 0.85 0.93
+Moderate reliability (pabak < 0.69)
+Policy 0.55 (77.7%) 0.526 0.78 0.89
+Politics 0.48 (74.1%) 0.425 0.79 0.92
+Average 0.71 (85.7%) 0.693 0.80 0.91
+Table 6: Human inter-annotator agreement scores (pabak and percentage), human-model agreement (pabak), and annotator confidence scores ([0-1]) across all nine topical frames introduced in Table 3.
+4.4.2.  Topical Frames
+Two expert annotators independently label a stratified random sample of 200 articles, providing nine binary labels with corresponding [0,1] confidence scores for each article.
+Significant class imbalance is a common challenge in multi-label annotation, and the inter-annotator agreement must consider the high base rate agreement where annotators agree on a label being absent from an article.
+The average agreement on label absence, over all nine topical frames, accounts for 65.82% of our evaluation sample, with the positive class (for any label) representing a minority of instances.
+Due to this prevalence bias, traditional measures like Cohen’s or Fleiss’ $\kappa$ are overly pessimistic reliability measures for our data.
+For a more accurate assessment that accounts for this imbalance, we report the prevalence-adjusted bias-adjusted kappa, or pabak Byrt et al.
+(1993), which is especially crucial for evaluation of consensus on the rarer labels.
+The inter-annotator agreement results, measured by pabak, are presented in Table 6 and demonstrate substantial agreement on average (0.71, with 85.7% agreement).
+Three topical frame categories — Welfare, Public Opinion, and Legal — reveal near-perfect agreement (>> 0.8).
+Overall, seven of the nine labels achieve pabak scores $\geq$ 0.7, indicating a consistent and shared understanding of the annotation framework.
+The two remaining classes, Policy and Politics, have moderate agreement, as they represent more abstract and often overlapping conceptual domains.
+Further, we examine the correlation between the average confidence scores reported by the human experts and those reported by phi-4.
+The framing categories are complex and sometimes overlapping, resulting in a weak positive correlation $\rho$ = 0.37.
+Notably, phi-4 exhibits systematically higher confidence (avg.
+0.91) than human annotators (avg.
+0.80) across all categories, regardless of the extent of agreement.
+This suggests that users should not interpret phi-4’s self-reported confidence scores as reliable indicators of annotation quality.
+Implications for Dataset Use: The strong agreement validates the overall reliability of our annotation schema. However, we recommend that users consider the following nuances in their studies:
+  • 1)
+    Frames with high/substantial reliability (pabak > 0.69) are suitable for fine-grained analysis such as tracking shifts in Welfare framing after major tragedies, or comparing Security or Legal discourse across state-aligned media.
+  • 2)
+    The moderately reliable frames (pabak < 0.69) — Policy and Politics — yield lower agreement due to their inherent conceptual overlap.
+These frame labels may be best suited for coarse-grained topic modeling or as supplementary features, with the understanding that they represent broader discourse categories that often intersect.
+As such, their unification may provide more reliable signals.
+  • 3)
+    The expert-annotated sample of 200 documents, stratified across sources, languages, and time periods, provides a completely validated high-quality gold-standard benchmark.
+Our stratified sampling ensures it is suitable for evaluating the performance of topical frame classification systems on real-world data on contentious wartime narratives.
+Together, these evaluations demonstrate that dnipro’s annotations achieve a balance of scale and quality rare in wartime media corpora: automated methods enable comprehensive coverage, while stratified human validation ensures researchers can make informed decisions regarding labels, as per their task-specific goals.
+The annotation guidelines and task-specific prompts are provided in Appendix A and C, respectively.
+5.  Use Case Experiments
+To demonstrate the analytical potential of dnipro, we present four use-case experiments exploiting its unique multi-perspective structure.
+These are not exhaustive analyses, but proof-of-concept demonstrations showing how the corpus enables diverse research in computational journalism, political communication, discourse analysis, and information warfare studies.
 Each application highlights specific affordances of the corpus while suggesting broader research directions.
-Thematic Analysis of Conflict Events:
-Our analyses focus on four major conflict events that received substantial global coverage and generated clear geopolitical divisions, inviting transnational comparisons (see Table 4).
-For each event, we filter articles using specific location identifiers (e.g., Bucha) to ensure event-specific analysis.
-To examine how media outlets prioritize different themes when covering conflict events, we analyze the topical content and emphasis patterns across sources.
-We identify which concepts, actors, and subtopics receive attention in coverage of major events, revealing how outlets differentially emphasize aspects such as historical context, humanitarian concerns, or institutional responses.
-Using GPT-3.5 Ye et al.
-(2023) with structured prompts, we tag articles from weekly time windows surrounding each major event to identify key concepts, actors, and subtopics.
-For the events described in Table 4, we obtain 810 articles, with 10.5 such tags (avg.) per article.
-We aggregate these tags by event and source country to find the distinct thematic emphasis patterns across media outlets.
-Russian and Chinese outlets consistently emphasize institutional terminology (special military operation), security concerns (nuclear weapons), and counter-narratives (media manipulation, ceasefire), reflecting strategic adversarial framing.
-Ukrainian sources predominantly highlight humanitarian dimensions (civilian casualties) with frequent references to specific conflict locations.
-American and British media balance humanitarian topics with institutional responses (sanctions, international criminal court) and policy discussions (no fly zone).
-These patterns demonstrate how dnipro captures differences in thematic prioritization across geopolitical actors even when covering the same underlying events.
-Russian Media Ukrainian Media
-Russia had testimony from residents of Bucha and Mariupol, who reported brutal crimes, such as torture, committed by Ukrainian nationalists.
-(TASS) All eyewitnesses noted the systematic nature of the killings in Bucha by the Russian army.
-(Censor)
-Russian forces jointly with border guards repelled these attacks and hit enemy positions near the border and in the Sumy Region.
-(TASS) Russian forces attacked six border areas and settlements of Sumy Oblast on June 7, firing 14 times and causing at least 54 explosions, the Sumy Oblast Military Administration reported.
-(Kyiv Independent)
-Table 7: Contradictory claims from opposing media outlets, highlighting how conflicting narratives are constructed not through direct refutation, but via divergent attribution, framing, and selection biases.
-Topical Framing:
-We conduct a corpus-wide analysis using our frame definitions (Table 3).
+Figure 2: Topical framing across news outlets (for each frame, % of articles with confidence > 0.5): Security dominates all outlets, while Economy is most prominent in Chinese media and Financial Times (ft).
+Other abbr.: Kyiv Independent (ki), Sputnik (sp), European Pravda (ep), Global Times (gt), China Daily (cd), Xinhua (xh).
+1.
+A Thematic Study of Conflict Events: Our analysis focuses on the four major conflict events shown in Table 4, which received substantial global coverage, generated clear geopolitical divisions, and invited transnational comparisons.
+For each event, we filter articles using specific location identifiers (e.g., Bucha) to ensure relevance.
+We examine how conflict events are framed by analyzing patterns of thematic salience across news outlets.
+By identifying which concepts, actors, and subtopics are underscored in coverage, we reveal how outlets differentially emphasize aspects such as historical context, humanitarian concerns, or institutional responses.
+Word clouds visualizing these thematic patterns for each conflict event and outlet country are provided in Appendix B.
+Our prelimnary analysis shows that Russian and Chinese outlets consistently emphasize institutional terminology (special military operation), security concerns (nuclear weapons), and counter-narratives (media manipulation, ceasefire), reflecting strategic adversarial framing.
+Ukrainian outlets predominantly highlight humanitarian dimensions (civilian casualties) with frequent references to specific conflict locations.
+Media from the USA and the UK balance humanitarian topics with institutional responses (sanctions, international criminal court) and policy discussions (no fly zone).
+Thus, dnipro captures differences in thematic salience and agenda-setting across geopolitical actors narrating the same underlying events.
+2.
+Topical Framing: We conduct a corpus-wide analysis using the schema introduced in Table 3.
 Article-level topical frames provide a comparable map of what each media outlet emphasizes, revealing systematic priorities that contextualize downstream analyses such as stance, sentiment, or discourse.
-Articles are annotated with phi-4 using a few-shot instruction prompt, producing multi-label assignments along with confidence scores.
-Figure 2 shows that Security dominates across all sources, with the highest shares in Ukrainian media, followed by Russian state media and Western sources like CNN.
-The second-most prevalent category is Politics, elevated in Russian and Chinese sources.
-It is substantially present in other sources too.
-Economy is most visible in Chinese media and, naturally, The Financial Times, while Policy appears steadily across sources, with higher emphasis in Chinese and British coverage.
+Articles are annotated with phi-4 using a few-shot instruction prompt, producing multi-label assignments along with confidence scores (as described in § 4.4.2).
+Figure 2 shows that Security dominates across all outlets, with the highest shares in Ukrainian media, followed by Russian state media and Western outlets like CNN.
+The second-most prevalent category is Politics, elevated in Russian and Chinese media.
+It is substantially present in other outlets too.
+Economy is most visible in Chinese media and, naturally, The Financial Times, while Policy appears steadily across outlets, with higher emphasis in Chinese media and the UK.
 Relative to others, Russian media leans more on Legal topics.
 Notably, discussions of Morality are extremely rare despite the focal topic of this dataset being an extended war, indicating that operational, strategic, and material considerations far outweigh explicit ethical rhetoric in war coverage.
-Figure 3: Daily average of stance scores towards Ukraine, grouped by media sources from five nations.
-The first 15 days of the war (left), contrasted with the last 15 days collected for dnipro (right).
-Stance analysis:
-Based on our stance annotations in § 4.3, we provide a comparison of the stance towards Ukraine between the first fortnight of the war and the last 15 days of dnipro data.
-This reveals a striking temporal shift across geopolitical lines.
-In Figure 3, we see that Chinese media was slightly pro-Ukrainian.
-By August 2024, however, this shifted to mild opposition.
-Russian coverage is consistently negative, but there is a drastic negative shift.
-The US media stance becomes even more supportive, while UK news begins to exhibit greater variability.
-This temporal comparison demonstrates how dnipro can provide analytical insights into reporting patterns.
-In this use case, it illustrates the polarization of geopolitical alignments.
-Divergent Narratives:
-To investigate how media outlets construct and contest narratives around specific flash points in the Russo-Ukrainian war, we analyze claims to identify conflicting representations of key events.
+3.
+Stance Analysis: We compare stance toward Ukraine between the war’s first 15 days and the final 15 days of data collection (Figure 3), revealing striking temporal shifts.
+Chinese media shifted from mildly pro-Ukrainian to mildly oppositional, likely reflecting China’s deepening strategic alignment with Russia.
+Russian coverage moved from a mildly negative stance to overwhelmingly hostile – possibly correlating with battlefield changes and political rhetoric.
+The US coverage became even more supportive, while the UK exhibited greater variability, possibly due to the transition from early solidarity to war fatigue, domestic economic pressures and/or leadership changes.
+The temporal dimension of dnipro proves crucial for these insights: China’s pivot, Russian and American intensification in opposite directions, and growing ambivalence in the UK.
+Figure 3: Daily average stance scores toward Ukraine, grouped by nation, comparing the first 15 days of the war (left) with the last 15 days of data collection (right).
+Note the dramatic intensification of Russian negativity and American support, increased variability in the UK, and the gentle negative shift in Chinese media.
+4.
+Divergent Narratives: To investigate how media outlets construct and contest narratives Gamson et al.
+(1992); Takeshita (1997), we analyze claims to identify conflicting representations of key events.
 Building on prior work in conflict journalism Entman (2010); Thussu et al.
 (2017), this analysis unearths how opposing perspectives are embedded in individual claims.
-Using weekly data subsets for each event, we extract sentence-level claims and categorize them into six types adapted from prior factuality and attribution typologies (Kotonya and Toni, 2020; Spinde et al., 2021):
+Using weekly data for each event, we extract sentence-level claims and categorize them into six types, motivated by epistemology Chafe and Nichols (1986), subjectivity and stance analysis Wiebe et al.
+(2005); Ferreira and Vlachos (2016), and speech acts Austin (1962); Searle (1969):
+  • factual: direct statements of observable or confirmed events;
+  • attributed: claims attributed to specific actors;
+  • disputed: explicit denials or rebuttals;
+  • reported: indirect accounts from third-parties;
+  • command/policy: official statements or directives;
+  • analytical/assessment: interpretive or evaluative judgments by experts or institutions.
+With GPT-3.5 plus human-in-the-loop validation, we label claims via structured prompts, and group semantically related claims through agglomerative clustering on Sentence Transformer representations Reimers and Gurevych (2019).
+Our key findings demonstrate dnipro’s utility in the study of information warfare, perception, and distortion Galtung and Ruge (1965) — not as explicit propaganda, but as systematic divergence of constructed realities across geopolitical divides:
+  • Contradictions are usually implicit: Contradictions predominantly emerge across geopolitical divides as competing interpretations.
+Outlets rarely offer explicit rebuttals, instead they construct incompatible narratives through different topical frames, omissions, or counter-attributions and selective emphasis (Table 7).
+  • Alignment with geopolitical blocs: Entailment occurs primarily within aligned media ecosystems: Ukrainian and Western outlets reinforce each other through overlapping factual, reported, or command/policy claims, while Russian media maintains internal consistency through attributed and analytical claims citing official statements.
+  • Framing substitutes for rebuttal: When covering the same event (e.g., an airstrike), Russian media might characterize it as precision targeting of military assets (factual/attributed), while Western media offer contrary eyewitness accounts (disputed/reported).
+Bucha Massacre
+ru Russia had testimony from residents of Bucha and Mariupol, who reported brutal crimes, such as torture, committed by Ukrainian nationalists. (TASS)
+ua All eyewitnesses noted the systematic nature of the killings in Bucha by the Russian army. (Censor)
+Sumy Oblast Attacks
+ru Russian forces jointly with border guards repelled these attacks and hit enemy positions near the border and in the Sumy Region. (TASS)
+ua Russian forces attacked six border areas and settlements of Sumy Oblast on June 7, firing 14 times and causing at least 54 explosions …(Kyiv Independent)
+Table 7: Competing realities narrated by Russian and Ukrainian media: Russian sources attribute atrocities to Ukrainian forces (yellow), while Ukrainian sources blame Russian forces (red).
+Narratives are opposed in attribution and framing, absent direct refutation or negation.
+6.  Conclusion
+We present dnipro, a longitudinal corpus of 246K news articles capturing competing geopolitical narratives of the Russo-Ukrainian war across 31 months (Feb 2022 – Aug 2024).
+Unlike prior corpora focused on social media volume or single-language coverage, dnipro enables comparative analysis across nation-states through structured, multilingual news data with human-evaluated annotations for stance, sentiment, and topical framing.
+Our experiments demonstrate dnipro’s capacity to support diverse analytical approaches and reveal systematic patterns, from thematic emphasis patterns and temporal stance tracking to fine-grained claim-level contradiction analysis.
+This makes dnipro uniquely suited for studying information ecosystems where competing narratives coexist without direct engagement, and contradictions are constructed implicitly: media outlets build parallel realities through divergent attribution and selective emphasis rather than direct refutation.
+This has methodological implications: traditional approaches assume explicit falsehoods or rebuttals Scantamburlo and Zollo (2026); Guo et al.
+(2022), but competing narratives operate through incompatible framing.
+Thus, data-driven research needs new methods that map how attribution choices and temporal patterns construct divergent realities across information ecosystems.
+dnipro enables scalability in several research directions: (a) longitudinal narrative evolution (e.g., what are the predictors of geopolitical realignments); (b) cross-lingual information flow (e.g., which media outlets serve as bridges, and which ones serve as echo chambers); (c) event-driven polarization: (e.g., can we predict which events trigger lasting narrative shifts); and (d) computational detection of incompatible narrative frames, in the absence of explicit disagreement markers.
+As international conflicts unfold across the fragmented media landscape, understanding these dynamics becomes essential for researchers, policymakers, journalists, and citizens.
+dnipro is publicly available under CC-BY-NC 4.0:
+We invite researchers to utilize dnipro, and develop new methods for detecting, tracking, and understanding narrative divergence: work that bridges computational analysis with urgent questions about truth, persuasion, and power in the digital age.
+7.  Limitations
+Temporal scope: dnipro covers February 2022 through August 2024, but the Russo-Ukrainian War has not concluded.
+Thus, significant developments after this period — including shifts in foreign policy, territorial control, or international alliances — are not captured.
+Researchers studying more recent events should consider this temporal boundary when interpreting findings, or extend our corpus to include later developments.
+Geographic coverage: Our selection stresses on geopolitical actors directly involved (Russia and Ukraine) and influential despite no direct engagement in the war (the United States, the United Kingdom, and China).
+NATO representation, however, is limited to two of its thirty-two member states.
+Perspectives from several influential NATO states (e.g., Germany and its state-own media, Deutsche Welle) are absent.
+Regional actors and nations with complex positions may also be included in the analyses we have underscored in our work.
+Such actors include Hungary and Romania, both of which border Ukraine; Belarus, also bordering Ukraine, but currently allied with the Russian state; Türkiye, a powerful regional actor, especially in the Black Sea; Georgia, a NATO contender that continues to maintain reasonably strong relations with Russia; and India, which maintains nuanced multi-alignments.
+The inclusion of these actors may reveal more sophisticated patterns of alignment and hedging than the polarized perspectives currently represented.
+Annotation reliability: While our human evaluations demonstrate substantial inter-annotator agreement (§ 4.4), automated ground-truth annotations are never perfect.
+This is not a limitation for qualitative analyses, mixed-method approaches, or empirical analyses for coarse-grained signals; but research requiring finer labels for specific nuanced tasks may benefit from additional human validation of their task-relevant subsets of dnipro.
+Scope of use-case experiments: Our exploratory analyses demonstrate analytical possibilities rather than provide definitive empirical claims.
+These demonstrations are not limitations of the corpus.
+Rather, they serve as proof-of-concept for the types of research dnipro enables: longitudinal stance tracking, framing comparison across media outlets and nation states, and contradiction detection.
+But we emphasize that these are not exhaustive treatments of these topics.
+Each use case is illustrative, and intended for deeper investigations.
+Translation and linguistic nuances: Non-English articles are machine-translated to English for cross-lingual analysis.
+Even as this enables systematic comparisons, translations may lose culturally specific rhetoric or connotations that carry meaning in the original language.
+Researchers with relevant language expertise may benefit from analyzing the untranslated Russian and Chinese Mandarin articles for such discursive phenomena.
+8.  Ethical Considerations
+dnipro documents an ongoing conflict with significant human costs and geopolitical implications. We acknowledge several potential ethical dimensions:
+  • Dual-use considerations.
+This corpus enables research into how narratives shape perception, with applications in computational journalism, political communication, and information literacy.
+However, the same analytical methods could be misused to refine propaganda or manipulate opinion.
+We urge researchers to consider the societal implications of their work and to strive for transparency, cross-cultural understanding, and informed discourse.
+  • Source inclusion and representation.
+dnipro inherently reflects the biases present in global media coverage.
+The inclusion of opposing narratives—including state-controlled media—is a deliberate choice, to empower comparative analyses.
+Inclusion does not imply endorsement of content.
+Some sources may contain disinformation or inflammatory rhetoric.
+We include them because the research possibilitied outlined earlier require access to the full spectrum of discourse, not sanitized versions.
+Researchers should contextualize sources appropriately using the provided metadata (media outlet ownership, editorial stance, etc.).
+We strongly recommend any derived datasets or findings to retain all the provided metadata, to ensure that articles can be interpreted within their proper editorial and temporal contexts.
+  • Sensitive content.
+This corpus pertains to an armed conflict.
+Many articles may describe violence and other distressing aspects of armed conflicts.
+Researchers working with this data should be aware of this, particularly when conducting qualitative analysis or annotation tasks.
+We recommend appropriate support structures for research teams.
+  • Data provenance and access, All content was gathered from publicly available news sources without breaching access restrictions or terms of service.
+We provide URLs rather than full text to respect intellectual property rights and enable verification, and do not include any copyright-protected content or any private communications.
+The CC-BY-NC 4.0 license permits research and educational use while prohibiting commercial exploitation.
+We designed dnipro to support research that illuminates how information ecosystems function during contested events — work that ultimately serves democratic accountability, informed citizenship, and conflict resolution.
+We earnestly welcome community input on additional safeguards or ethical considerations.
+9.  Bibliographical References
+  • AFP (2022) Russia Bans Media Outlets From Using Words ‘War’, ‘Invasion’. Note: The Moscow TimesDate: Feb 6, 2022; Retrieved: Nov 29, 2024 External Links: Link Cited by: footnote 1.
+  • A. U. Akash, A. Fahmy, and A. Trabelsi (2025) Can Large Language Models Address Open-Target Stance Detection?. External Links: 2409.00222, Link Cited by: Appendix A. Annotation Methodology.
+  • H.
+M.
+Alemayehu, H.
+M.
+Zahera, and A.
+Ngonga Ngomo (2024) Error analysis of multilingual language models in machine translation: a case study of English-Amharic translation.
+In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Y.
+Al-Onaizan, M.
+Bansal, and Y.
+Chen (Eds.), Miami, Florida, USA, pp.
+19758–19768.
+External Links: Document Cited by: §4.2.
+  • M.
+Alyukov, M.
+Kunilovskaya, and A.
+Semenov (2023) Wartime Media Monitor (WarMM-2022): A Study of Information Manipulation on Russian Social Media during the Russia-Ukraine War.
+In Proceedings of the 7th Joint SIGHUM Workshop on Computational Linguistics for Cultural Heritage, Social Sciences, Humanities and Literature, S.
+Degaetano-Ortlieb, A.
+Kazantseva, N.
+Reiter, and S.
+Szpakowicz (Eds.), Dubrovnik, Croatia, pp.
+152–161.
+External Links: Document Cited by: §2.
+  • J. L. Austin (1962) How to Do Things with Words. Oxford University Press. Cited by: §5.
+  • A.
+Bawa, U.
+Kursuncu, D.
+Achilov, V.
+L.
+Shalin, N.
+Agarwal, and E.
+Akbas (2025) Telegram as a Battlefield: Kremlin-Related Communications During the Russia-Ukraine Conflict.
+Proceedings of the International AAAI Conference on Web and Social Media 19 (1), pp.
+2361–2370.
+External Links: Document Cited by: §2, §2.
+  • S.
+Bradshaw, M.
+Elswah, M.
+Haque, and D.
+Quelle (2024) Strategic Storytelling: Russian State-Backed Media Coverage of the Ukraine War.
+International Journal of Public Opinion Research 36 (3), pp.
+edae028.
+External Links: Document Cited by: §1.
+  • T.
+B.
+Brown, B.
+Mann, N.
+Ryder, M.
+Subbiah, J.
+Kaplan, P.
+Dhariwal, A.
+Neelakantan, P.
+Shyam, G.
+Sastry, A.
+Askell, S.
+Agarwal, A.
+Herbert-Voss, G.
+Krueger, T.
+Henighan, R.
+Child, A.
+Ramesh, D.
+M.
+Ziegler, J.
+Wu, C.
+Winter, C.
+Hesse, M.
+Chen, E.
+Sigler, M.
+Litwin, S.
+Gray, B.
+Chess, J.
+Clark, C.
+Berner, S.
+McCandlish, A.
+Radford, I.
+Sutskever, and D.
+Amodei (2020) Language models are few-shot learners.
+External Links: 2005.14165, Link Cited by: 3rd item.
+  • T. Byrt, J. Bishop, and J. B. Carlin (1993) Bias, prevalence and kappa. Journal of Clinical Epidemiology 46 (5), pp. 423–429. External Links: Document Cited by: §4.4.2.
+  • W. Chafe and J. Nichols (Eds.) (1986) Evidentiality: The Linguistic Coding of Epistemology. Advances in Discourse Processes, Vol. 20, Ablex, Norwood, NJ. Cited by: §5.
+  • E.
+Chen and E.
+Ferrara (2023) Tweets in Time of Conflict: A Public Dataset Tracking the Twitter Discourse on the War between Ukraine and Russia.
+Proceedings of the International AAAI Conference on Web and Social Media 17 (1), pp.
+1006–1013.
+External Links: Document Cited by: §2.
+  • L.
+Derviş (2023) Transformation of Geopolitical Perceptions in the Russian-Ukrainian War: Impact on Regional Relations in the Future.
+Futurity of Social Sciences 1 (1), pp.
+21–34.
+External Links: Document Cited by: §1.
+  • R. M. Entman (1993) Framing: Toward Clarification of a Fractured Paradigm. Journal of Communication 43 (4), pp. 51–58. External Links: Document Cited by: footnote 2.
+  • R. M. Entman (2010) Media framing biases and political power: Explaining slant in news of Campaign 2008. Journalism 11 (4), pp. 389–408. External Links: Document Cited by: §5.
+  • N. Fairclough (2010) Critical Discourse Analysis: The Critical Study of Language. 2 edition, Routledge. External Links: Document Cited by: footnote 2.
+  • A.
+Fan, S.
+Bhosale, H.
+Schwenk, Z.
+Ma, A.
+El-Kishky, S.
+Goyal, M.
+Baines, O.
+Celebi, G.
+Wenzek, V.
+Chaudhary, N.
+Goyal, T.
+Birch, V.
+Liptchinsky, S.
+Edunov, M.
+Auli, and A.
+Joulin (2021) Beyond English-Centric Multilingual Machine Translation.
+Journal of Machine Learning Research 22 (107), pp.
+1–48.
+External Links: Link Cited by: §4.2.
+  • W.
+Ferreira and A.
+Vlachos (2016) Emergent: a novel data-set for stance classification.
+In Proceedings of the 2016 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, K.
+Knight, A.
+Nenkova, and O.
+Rambow (Eds.), San Diego, California, pp.
+1163–1168.
+External Links: Document Cited by: §5.
+  • J.
+Galtung and M.
+H.
+Ruge (1965) The Structure of Foreign News: The Presentation of the Congo, Cuba and Cyprus Crises in Four Norwegian Newspapers.
+Journal of Peace Research 2 (1), pp.
+64–91.
+External Links: Document Cited by: §5.
+  • W.
+A.
+Gamson, D.
+Croteau, W.
+Hoynes, and T.
+Sasson (1992) Media Images and the Social Construction of Reality.
+Annual Review of Sociology 18 (1), pp.
+373–393.
+External Links: Document Cited by: §5.
+  • W.
+A.
+Gamson and A.
+Modigliani (1989) Media Discourse and Public Opinion on Nuclear Power: A Constructionist Approach.
+American Journal of Sociology 95 (1), pp.
+1–37.
+External Links: Document Cited by: footnote 2.
+  • Z.
+Guo, M.
+Schlichtkrull, and A.
+Vlachos (2022) A Survey on Automated Fact-Checking.
+Transactions of the Association for Computational Linguistics 10, pp.
+178–206.
+External Links: Document Cited by: §6.
+  • J. Habermas (1991) Erläuterungen zur Diskursethik. Suhrkamp Taschenbuch Wissenschaft, Suhrkamp, Frankfurt am Main. Cited by: item 3, Table 3.
+  • S.
+Hakimov and G.
+S.
+Cheema (2024) Unveiling Global Narratives: A Multilingual Twitter Dataset of News Media on the Russo-Ukrainian Conflict.
+In Proceedings of the 2024 International Conference on Multimedia Retrieval, ICMR ’24, New York, NY, USA, pp.
+1160–1164.
+External Links: Document Cited by: §2, §2.
+  • J.
+Hartmann, M.
+Heitmann, C.
+Siebert, and C.
+Schamp (2023) More than a Feeling: Accuracy and Application of Sentiment Analysis.
+International Journal of Research in Marketing 40 (1), pp.
+75–87.
+External Links: Document Cited by: §4.3, 2nd item, item (b).
+  • M. Honnibal, I. Montani, S. Van Landeghem, and A. Boyd (2020) spaCy: Industrial-strength Natural Language Processing in Python. External Links: Document Cited by: §4.3, item (a).
+  • M.
+Ibrahim, B.
+Wang, M.
+Xu, and H.
+Xu (2025) A multidimensional analysis of media framing in the Russia-Ukraine war.
+Journal of Computational Social Science 8 (34).
+External Links: Document Cited by: §2.
+  • E.
+Johais and M.
+Meis (2024) ‘Unleash the hounds!’: NAFO’s memetic war narrative on the Russo-Ukrainian conflict.
+Critical Studies on Security 13 (1), pp.
+1–13.
+External Links: Document Cited by: §1.
+  • N.
+Khairova, A.
+Galassi, F.
+L.
+Scudo, B.
+Ivasiuk, and I.
+Redozub (2024) Unsupervised approach for misinformation detection in Russia-Ukraine war news.
+In Proceedings of the 8th International Conference on Computational Linguistics and Intelligent Systems.
+Volume IV: Computational Linguistics Workshop, pp.
+21–36.
+External Links: Link Cited by: §2.
+  • Microsoft (2024) Phi-4 Technical Report. Note: Accessed: Oct 17, 2025 External Links: Link Cited by: §4.3, §4.3, 1st item, Appendix A. Annotation Methodology.
+  • P.
+Nemkova, S.
+Ubani, S.
+O.
+Polat, N.
+Kim, and R.
+D.
+Nielsen (2023) Detecting Human Rights Violations on Social Media during Russia-Ukraine War.
+External Links: 2306.05370, Document, Link Cited by: §2, §2.
+  • G.
+Nygren, M.
+Glowacki, J.
+Hök, I.
+Kiria, D.
+Orlova, and D.
+Taradai (2016) Journalism in the Crossfire: Media Coverage of the War in Ukraine in 2014.
+Journalism Studies 19 (7), pp.
+1059–1078.
+External Links: Document Cited by: §1.
+  • OpenAI (2023) GPT-3.5 language model. Note: https://developers.openai.com/api/docs/models/gpt-3.5-turbo Cited by: 3rd item, Thematic Study of Conflict Events:.
+  • B. Pancevski (2024) One Million Are Now Dead or Injured in the Russia-Ukraine War. Note: The Wall Street JournalDate: Sep 17, 2024; Retrieved: Feb 3, 2025 External Links: Link Cited by: §1.
+  • C.
+Y.
+Park, J.
+Mendelsohn, A.
+Field, and Y.
+Tsvetkov (2022) Challenges and Opportunities in Information Manipulation Detection: An Examination of Wartime Russian Media.
+In Findings of the Association for Computational Linguistics: EMNLP 2022, Y.
+Goldberg, Z.
+Kozareva, and Y.
+Zhang (Eds.), pp.
+5209–5235.
+External Links: Document Cited by: §2, §2.
+  • N.
+Reimers and I.
+Gurevych (2019) Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks.
+In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), K.
+Inui, J.
+Jiang, V.
+Ng, and X.
+Wan (Eds.), pp.
+3982–3992.
+External Links: Document Cited by: §5.
+  • N.
+Roman, W.
+Wanta, and I.
+Buniyak (2017) Information wars: Eastern Ukraine military conflict coverage in the Russian, Ukrainian and U.S.
+newscasts.
+International Communication Gazette 79 (4), pp.
+357–378.
+External Links: Document Cited by: §1.
+  • M.
+Saletta, A.
+Kruger, T.
+Primoratz, A.
+Barnett, T.
+van Gelder, and R.
+E.
+Horn (2020) The role of narrative in collaborative reasoning and intelligence analysis: A case study.
+PLoS One 15 (1), pp.
+e0226981.
+External Links: Document Cited by: §1.
+  • T. Scantamburlo and F. Zollo (2026) A Scoping Review of Misinformation Research: Definitions, Conceptualizations, and Measurement Approaches. ACM Trans. Web. External Links: Document Cited by: §6.
+  • J. R. Searle (1969) Speech Acts: An Essay in the Philosophy of Language. Cambridge University Press. External Links: Document Cited by: §5.
+  • A. Shevtsov, C. Tzagkarakis, D. Antonakaki, P. Pratikakis, and S. Ioannidis (2022) Twitter Dataset on the Russo-Ukrainian War. arXiv 2204.08530. External Links: Document Cited by: §2.
+  • J.
+Silvertsen (2024) Projecting Geopolitical Power and Acting as One: The EU’s Foreign Policy in the Russian-Ukrainian Conflict [2014 and 2022-23].
+Master’s Thesis, Norwegian University of Science and Technology, Trondheim, Norway.
+External Links: Link Cited by: §1.
+  • A.
+V.
+Smirnov, N.
+Teslya, N.
+Shilov, D.
+Frank, E.
+Minina, and M.
+Kovacs (2022) Quantitative Comparison of Translation by Transformers-Based Neural Network Models.
+In International Conference on Enterprise Information Systems, External Links: Document Cited by: §4.2.
+  • J.
+Susanne Pohl, S.
+Markmann, D.
+Assenmacher, and C.
+Grimme (2023) Invasion@Ukraine: Providing and Describing a Twitter Streaming Dataset That Captures the Outbreak of War between Russia and Ukraine in 2022.
+Proceedings of the International AAAI Conference on Web and Social Media 17 (1), pp.
+1093–1101.
+External Links: Document Cited by: §2, §2.
+  • T.
+Takeshita (1997) Exploring the Media’s Roles in Defining Reality: From Issue-Agenda Setting to Attribute-Agenda Setting.
+In Communication and Democracy: Exploring the Intellectual Frontiers in Agenda-Setting Theory, M.
+McCombs, D.
+L.
+Shaw, and D.
+Weaver (Eds.), pp.
+15–27.
+External Links: Link Cited by: §5.
+  • N. Team (2022) No Language Left Behind: Scaling Human-Centered Machine Translation. External Links: 2207.04672, Link Cited by: §4.2.
+  • D. K. Thussu, H. De Burgh, and A. Shi (Eds.) (2017) China’s Media Go Global. Routledge, London. External Links: Document Cited by: §1, §5.
+  • J.
+Tiedemann, M.
+Aulamo, D.
+Bakshandaeva, M.
+Boggia, S.
+Grönroos, T.
+Nieminen, A.
+Raganato, Y.
+Scherrer, R.
+Vazquez, and S.
+Virpioja (2023) Democratizing Neural Machine Translation with OPUS–MT.
+Language Resources and Evaluation 58, pp.
+713–755.
+External Links: Document Cited by: §4.2.
+  • J.
+Watling, O.
+V.
+Danylyuk, and N.
+Reynolds (2023) Preliminary Lessons from Russia’s Unconventional Operations During the Russo-Ukrainian War, February 2022–February 2023.
+Note: Royal United Services Institute (RUSI) for Defence and Security StudiesDate: Mar 29, 2023; Retrieved: Nov 28, 2024 External Links: Link Cited by: §1.
+  • P. G. d. Werd (2018) Critical Intelligence: Analysis by Contrasting Narratives. Doctoral Thesis, Universiteit Utrecht. External Links: Link Cited by: §1.
+  • J.
+Wiebe, T.
+Wilson, and C.
+Cardie (2005) Annotating Expressions of Opinions and Emotions in Language.
+In Language Resources and Evaluation, Vol.
+39, pp.
+165–210.
+External Links: Document Cited by: §5.
+  • Q.
+Yu and A.
+Fliethmann (2022) Frame Detection in German Political Discourses: How Far Can We Go Without Large-Scale Manual Corpus Annotation?.
+Journal for Language Technology and Computational Linguistics 35 (2), pp.
+15–31.
+External Links: Document Cited by: Table 3.
+  • X.
+Zhang, H.
+Wasserman, and W.
+Mano (2016) China’s Expanding Influence in Africa: Projection, Perception and Prospects in Southern African Countries.
+Communicatio 42 (1), pp.
+1–22.
+External Links: Document Cited by: §1.
+  • Y. Zhu, E. Haq, L. Lee, G. Tyson, and P. Hui (2022) A Reddit Dataset for the Russo-Ukrainian Conflict in 2022. arXiv 2206.05107. External Links: Document Cited by: §2.
+10.  Language Resource References
+  • AFP (2022) Russia Bans Media Outlets From Using Words ‘War’, ‘Invasion’. Note: The Moscow TimesDate: Feb 6, 2022; Retrieved: Nov 29, 2024 External Links: Link Cited by: footnote 1.
+  • A. U. Akash, A. Fahmy, and A. Trabelsi (2025) Can Large Language Models Address Open-Target Stance Detection?. External Links: 2409.00222, Link Cited by: Appendix A. Annotation Methodology.
+  • H.
+M.
+Alemayehu, H.
+M.
+Zahera, and A.
+Ngonga Ngomo (2024) Error analysis of multilingual language models in machine translation: a case study of English-Amharic translation.
+In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Y.
+Al-Onaizan, M.
+Bansal, and Y.
+Chen (Eds.), Miami, Florida, USA, pp.
+19758–19768.
+External Links: Document Cited by: §4.2.
+  • M.
+Alyukov, M.
+Kunilovskaya, and A.
+Semenov (2023) Wartime Media Monitor (WarMM-2022): A Study of Information Manipulation on Russian Social Media during the Russia-Ukraine War.
+In Proceedings of the 7th Joint SIGHUM Workshop on Computational Linguistics for Cultural Heritage, Social Sciences, Humanities and Literature, S.
+Degaetano-Ortlieb, A.
+Kazantseva, N.
+Reiter, and S.
+Szpakowicz (Eds.), Dubrovnik, Croatia, pp.
+152–161.
+External Links: Document Cited by: §2.
+  • J. L. Austin (1962) How to Do Things with Words. Oxford University Press. Cited by: §5.
+  • A.
+Bawa, U.
+Kursuncu, D.
+Achilov, V.
+L.
+Shalin, N.
+Agarwal, and E.
+Akbas (2025) Telegram as a Battlefield: Kremlin-Related Communications During the Russia-Ukraine Conflict.
+Proceedings of the International AAAI Conference on Web and Social Media 19 (1), pp.
+2361–2370.
+External Links: Document Cited by: §2, §2.
+  • S.
+Bradshaw, M.
+Elswah, M.
+Haque, and D.
+Quelle (2024) Strategic Storytelling: Russian State-Backed Media Coverage of the Ukraine War.
+International Journal of Public Opinion Research 36 (3), pp.
+edae028.
+External Links: Document Cited by: §1.
+  • T.
+B.
+Brown, B.
+Mann, N.
+Ryder, M.
+Subbiah, J.
+Kaplan, P.
+Dhariwal, A.
+Neelakantan, P.
+Shyam, G.
+Sastry, A.
+Askell, S.
+Agarwal, A.
+Herbert-Voss, G.
+Krueger, T.
+Henighan, R.
+Child, A.
+Ramesh, D.
+M.
+Ziegler, J.
+Wu, C.
+Winter, C.
+Hesse, M.
+Chen, E.
+Sigler, M.
+Litwin, S.
+Gray, B.
+Chess, J.
+Clark, C.
+Berner, S.
+McCandlish, A.
+Radford, I.
+Sutskever, and D.
+Amodei (2020) Language models are few-shot learners.
+External Links: 2005.14165, Link Cited by: 3rd item.
+  • T. Byrt, J. Bishop, and J. B. Carlin (1993) Bias, prevalence and kappa. Journal of Clinical Epidemiology 46 (5), pp. 423–429. External Links: Document Cited by: §4.4.2.
+  • W. Chafe and J. Nichols (Eds.) (1986) Evidentiality: The Linguistic Coding of Epistemology. Advances in Discourse Processes, Vol. 20, Ablex, Norwood, NJ. Cited by: §5.
+  • E.
+Chen and E.
+Ferrara (2023) Tweets in Time of Conflict: A Public Dataset Tracking the Twitter Discourse on the War between Ukraine and Russia.
+Proceedings of the International AAAI Conference on Web and Social Media 17 (1), pp.
+1006–1013.
+External Links: Document Cited by: §2.
+  • L.
+Derviş (2023) Transformation of Geopolitical Perceptions in the Russian-Ukrainian War: Impact on Regional Relations in the Future.
+Futurity of Social Sciences 1 (1), pp.
+21–34.
+External Links: Document Cited by: §1.
+  • R. M. Entman (1993) Framing: Toward Clarification of a Fractured Paradigm. Journal of Communication 43 (4), pp. 51–58. External Links: Document Cited by: footnote 2.
+  • R. M. Entman (2010) Media framing biases and political power: Explaining slant in news of Campaign 2008. Journalism 11 (4), pp. 389–408. External Links: Document Cited by: §5.
+  • N. Fairclough (2010) Critical Discourse Analysis: The Critical Study of Language. 2 edition, Routledge. External Links: Document Cited by: footnote 2.
+  • A.
+Fan, S.
+Bhosale, H.
+Schwenk, Z.
+Ma, A.
+El-Kishky, S.
+Goyal, M.
+Baines, O.
+Celebi, G.
+Wenzek, V.
+Chaudhary, N.
+Goyal, T.
+Birch, V.
+Liptchinsky, S.
+Edunov, M.
+Auli, and A.
+Joulin (2021) Beyond English-Centric Multilingual Machine Translation.
+Journal of Machine Learning Research 22 (107), pp.
+1–48.
+External Links: Link Cited by: §4.2.
+  • W.
+Ferreira and A.
+Vlachos (2016) Emergent: a novel data-set for stance classification.
+In Proceedings of the 2016 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, K.
+Knight, A.
+Nenkova, and O.
+Rambow (Eds.), San Diego, California, pp.
+1163–1168.
+External Links: Document Cited by: §5.
+  • J.
+Galtung and M.
+H.
+Ruge (1965) The Structure of Foreign News: The Presentation of the Congo, Cuba and Cyprus Crises in Four Norwegian Newspapers.
+Journal of Peace Research 2 (1), pp.
+64–91.
+External Links: Document Cited by: §5.
+  • W.
+A.
+Gamson, D.
+Croteau, W.
+Hoynes, and T.
+Sasson (1992) Media Images and the Social Construction of Reality.
+Annual Review of Sociology 18 (1), pp.
+373–393.
+External Links: Document Cited by: §5.
+  • W.
+A.
+Gamson and A.
+Modigliani (1989) Media Discourse and Public Opinion on Nuclear Power: A Constructionist Approach.
+American Journal of Sociology 95 (1), pp.
+1–37.
+External Links: Document Cited by: footnote 2.
+  • Z.
+Guo, M.
+Schlichtkrull, and A.
+Vlachos (2022) A Survey on Automated Fact-Checking.
+Transactions of the Association for Computational Linguistics 10, pp.
+178–206.
+External Links: Document Cited by: §6.
+  • J. Habermas (1991) Erläuterungen zur Diskursethik. Suhrkamp Taschenbuch Wissenschaft, Suhrkamp, Frankfurt am Main. Cited by: item 3, Table 3.
+  • S.
+Hakimov and G.
+S.
+Cheema (2024) Unveiling Global Narratives: A Multilingual Twitter Dataset of News Media on the Russo-Ukrainian Conflict.
+In Proceedings of the 2024 International Conference on Multimedia Retrieval, ICMR ’24, New York, NY, USA, pp.
+1160–1164.
+External Links: Document Cited by: §2, §2.
+  • J.
+Hartmann, M.
+Heitmann, C.
+Siebert, and C.
+Schamp (2023) More than a Feeling: Accuracy and Application of Sentiment Analysis.
+International Journal of Research in Marketing 40 (1), pp.
+75–87.
+External Links: Document Cited by: §4.3, 2nd item, item (b).
+  • M. Honnibal, I. Montani, S. Van Landeghem, and A. Boyd (2020) spaCy: Industrial-strength Natural Language Processing in Python. External Links: Document Cited by: §4.3, item (a).
+  • M.
+Ibrahim, B.
+Wang, M.
+Xu, and H.
+Xu (2025) A multidimensional analysis of media framing in the Russia-Ukraine war.
+Journal of Computational Social Science 8 (34).
+External Links: Document Cited by: §2.
+  • E.
+Johais and M.
+Meis (2024) ‘Unleash the hounds!’: NAFO’s memetic war narrative on the Russo-Ukrainian conflict.
+Critical Studies on Security 13 (1), pp.
+1–13.
+External Links: Document Cited by: §1.
+  • N.
+Khairova, A.
+Galassi, F.
+L.
+Scudo, B.
+Ivasiuk, and I.
+Redozub (2024) Unsupervised approach for misinformation detection in Russia-Ukraine war news.
+In Proceedings of the 8th International Conference on Computational Linguistics and Intelligent Systems.
+Volume IV: Computational Linguistics Workshop, pp.
+21–36.
+External Links: Link Cited by: §2.
+  • Microsoft (2024) Phi-4 Technical Report. Note: Accessed: Oct 17, 2025 External Links: Link Cited by: §4.3, §4.3, 1st item, Appendix A. Annotation Methodology.
+  • P.
+Nemkova, S.
+Ubani, S.
+O.
+Polat, N.
+Kim, and R.
+D.
+Nielsen (2023) Detecting Human Rights Violations on Social Media during Russia-Ukraine War.
+External Links: 2306.05370, Document, Link Cited by: §2, §2.
+  • G.
+Nygren, M.
+Glowacki, J.
+Hök, I.
+Kiria, D.
+Orlova, and D.
+Taradai (2016) Journalism in the Crossfire: Media Coverage of the War in Ukraine in 2014.
+Journalism Studies 19 (7), pp.
+1059–1078.
+External Links: Document Cited by: §1.
+  • OpenAI (2023) GPT-3.5 language model. Note: https://developers.openai.com/api/docs/models/gpt-3.5-turbo Cited by: 3rd item, Thematic Study of Conflict Events:.
+  • B. Pancevski (2024) One Million Are Now Dead or Injured in the Russia-Ukraine War. Note: The Wall Street JournalDate: Sep 17, 2024; Retrieved: Feb 3, 2025 External Links: Link Cited by: §1.
+  • C.
+Y.
+Park, J.
+Mendelsohn, A.
+Field, and Y.
+Tsvetkov (2022) Challenges and Opportunities in Information Manipulation Detection: An Examination of Wartime Russian Media.
+In Findings of the Association for Computational Linguistics: EMNLP 2022, Y.
+Goldberg, Z.
+Kozareva, and Y.
+Zhang (Eds.), pp.
+5209–5235.
+External Links: Document Cited by: §2, §2.
+  • N.
+Reimers and I.
+Gurevych (2019) Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks.
+In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), K.
+Inui, J.
+Jiang, V.
+Ng, and X.
+Wan (Eds.), pp.
+3982–3992.
+External Links: Document Cited by: §5.
+  • N.
+Roman, W.
+Wanta, and I.
+Buniyak (2017) Information wars: Eastern Ukraine military conflict coverage in the Russian, Ukrainian and U.S.
+newscasts.
+International Communication Gazette 79 (4), pp.
+357–378.
+External Links: Document Cited by: §1.
+  • M.
+Saletta, A.
+Kruger, T.
+Primoratz, A.
+Barnett, T.
+van Gelder, and R.
+E.
+Horn (2020) The role of narrative in collaborative reasoning and intelligence analysis: A case study.
+PLoS One 15 (1), pp.
+e0226981.
+External Links: Document Cited by: §1.
+  • T. Scantamburlo and F. Zollo (2026) A Scoping Review of Misinformation Research: Definitions, Conceptualizations, and Measurement Approaches. ACM Trans. Web. External Links: Document Cited by: §6.
+  • J. R. Searle (1969) Speech Acts: An Essay in the Philosophy of Language. Cambridge University Press. External Links: Document Cited by: §5.
+  • A. Shevtsov, C. Tzagkarakis, D. Antonakaki, P. Pratikakis, and S. Ioannidis (2022) Twitter Dataset on the Russo-Ukrainian War. arXiv 2204.08530. External Links: Document Cited by: §2.
+  • J.
+Silvertsen (2024) Projecting Geopolitical Power and Acting as One: The EU’s Foreign Policy in the Russian-Ukrainian Conflict [2014 and 2022-23].
+Master’s Thesis, Norwegian University of Science and Technology, Trondheim, Norway.
+External Links: Link Cited by: §1.
+  • A.
+V.
+Smirnov, N.
+Teslya, N.
+Shilov, D.
+Frank, E.
+Minina, and M.
+Kovacs (2022) Quantitative Comparison of Translation by Transformers-Based Neural Network Models.
+In International Conference on Enterprise Information Systems, External Links: Document Cited by: §4.2.
+  • J.
+Susanne Pohl, S.
+Markmann, D.
+Assenmacher, and C.
+Grimme (2023) Invasion@Ukraine: Providing and Describing a Twitter Streaming Dataset That Captures the Outbreak of War between Russia and Ukraine in 2022.
+Proceedings of the International AAAI Conference on Web and Social Media 17 (1), pp.
+1093–1101.
+External Links: Document Cited by: §2, §2.
+  • T.
+Takeshita (1997) Exploring the Media’s Roles in Defining Reality: From Issue-Agenda Setting to Attribute-Agenda Setting.
+In Communication and Democracy: Exploring the Intellectual Frontiers in Agenda-Setting Theory, M.
+McCombs, D.
+L.
+Shaw, and D.
+Weaver (Eds.), pp.
+15–27.
+External Links: Link Cited by: §5.
+  • N. Team (2022) No Language Left Behind: Scaling Human-Centered Machine Translation. External Links: 2207.04672, Link Cited by: §4.2.
+  • D. K. Thussu, H. De Burgh, and A. Shi (Eds.) (2017) China’s Media Go Global. Routledge, London. External Links: Document Cited by: §1, §5.
+  • J.
+Tiedemann, M.
+Aulamo, D.
+Bakshandaeva, M.
+Boggia, S.
+Grönroos, T.
+Nieminen, A.
+Raganato, Y.
+Scherrer, R.
+Vazquez, and S.
+Virpioja (2023) Democratizing Neural Machine Translation with OPUS–MT.
+Language Resources and Evaluation 58, pp.
+713–755.
+External Links: Document Cited by: §4.2.
+  • J.
+Watling, O.
+V.
+Danylyuk, and N.
+Reynolds (2023) Preliminary Lessons from Russia’s Unconventional Operations During the Russo-Ukrainian War, February 2022–February 2023.
+Note: Royal United Services Institute (RUSI) for Defence and Security StudiesDate: Mar 29, 2023; Retrieved: Nov 28, 2024 External Links: Link Cited by: §1.
+  • P. G. d. Werd (2018) Critical Intelligence: Analysis by Contrasting Narratives. Doctoral Thesis, Universiteit Utrecht. External Links: Link Cited by: §1.
+  • J.
+Wiebe, T.
+Wilson, and C.
+Cardie (2005) Annotating Expressions of Opinions and Emotions in Language.
+In Language Resources and Evaluation, Vol.
+39, pp.
+165–210.
+External Links: Document Cited by: §5.
+  • Q.
+Yu and A.
+Fliethmann (2022) Frame Detection in German Political Discourses: How Far Can We Go Without Large-Scale Manual Corpus Annotation?.
+Journal for Language Technology and Computational Linguistics 35 (2), pp.
+15–31.
+External Links: Document Cited by: Table 3.
+  • X.
+Zhang, H.
+Wasserman, and W.
+Mano (2016) China’s Expanding Influence in Africa: Projection, Perception and Prospects in Southern African Countries.
+Communicatio 42 (1), pp.
+1–22.
+External Links: Document Cited by: §1.
+  • Y. Zhu, E. Haq, L. Lee, G. Tyson, and P. Hui (2022) A Reddit Dataset for the Russo-Ukrainian Conflict in 2022. arXiv 2206.05107. External Links: Document Cited by: §2.
+Appendix A. Annotation Methodology
+This appendix provides complete prompt specifications and implementation details for the large-scale automated annotations described in § 4.3.
+All prompts follow structured output formats to enable reliable parsing and post-processing.
+Models and APIs: We used the following language models for automated annotations:
+  • microsoft/phi-4 (14B parameters) Microsoft (2024) for stance detection and topical framing;
+  • a fine-tuned RoBERTa-large model by Hartmann et al. (2023)666siebert/sentiment-roberta-large-english (Hugging Face). for sentiment analysis; and
+  • GPT-3.5 Turbo, via the OpenAI API (OpenAI (2023); Brown et al. (2020)), for thematic tagging and claim typology.
+Stance detection proceeds in two stages: (1) target generationidentifies stance-worthy entities or events in a sentence, and then (2) stance classificationdetermines the expressed attitude toward each target.
+This decomposition improves accuracy by separating entity recognition from sentiment interpretation Akash et al.
+(2025).
+We apply stance detection to articles from: (a) the first and last 15 days of data collection, and (b) four major conflict events listed in Table 4: the Bucha massacre, the Kursk offensive, Mariupol hospital attack, and Zaporizhzhia nuclear crisis.
+Target Generation.
+The model identifies all stance-worthy targets – actors, organizations, or events that can be supported or opposed – mentioned in a sentence.
+The prompt template for target generation is shown next:
+Stance Classification.
+For each detected target, the model classifies the stance expressed in the sentence using the surrounding passage as context, thereby resolving pronouns and implicit references.
+Sentiment analysis is performed at sentence-level, for key entities like Russia, Ukraine, Zelensky, Bucha, Kursk, etc.
+For each entity, we (a) extract all sentences mentioning the entity using the English-language model of spaCy Honnibal et al.
+(2020); (b) pass each sentence to the fine-tuned RoBERTa-large model Hartmann et al.
+(2023); (c) map model outputs to continuous scores in [-1, 1]; and (d) aggregate the scores daily and monthly, by source and country.
+This simple format provides sufficient context for the pretrained model, without requiring extensive prompt engineering.
+Topical framing is done using the nine categories described in the schema presented in Table 3.
+We perform this multiclass classification using phi-4 Microsoft (2024).
+The model receives an article’s title, subtitle (if present), and the full text content.
+As output, it then provides applicable frames along with confidence scores.
+Thematic tagging (see § 5, use case 1: “A Thematic Study of Conflict Events”) is done by extracting article-level tags using GPT-3.5 Turbo with one-shot prompting.
+The model is instructed to extract a set of specific tags summarizing each article’s content.
+The tags were intended to capture salient persons or people, locations, military units, political or narrative themes, and ideological frames surrounding major conflict events.
+To ensure a consistent structure that can support post-processing, the prompt was instructed to produce outputs in a strict JSON format.
+Claim typology was established with the six types of claims (see § 5, use case 4: “Divergent Narratives”): factual, attributed, disputed, reported, command/policy, and analytical/assessment.
+We use a structured two-shot prompting strategy with GPT-3.5 Turbo to extract and classify claims.
+Appendix B. Use Case Experiments
+Thematic Study of Conflict Events:
+We present word clouds in Figure 4 to visually support our thematic analysis (described in §  5.
+Using GPT-3.5 OpenAI (2023) with structured prompts, we tag articles from weekly time windows surrounding each major event to identify key concepts, actors, and subtopics.
+For the events described in Table 4, we obtain 810 articles, with 10.5 such tags (avg.) per article.
+We aggregate these tags by event and country to find the distinct thematic emphasis patterns across media outlets.
+In summary, each word cloud aggregates GPT-3.5-tagged concepts, actors, and subtopics by outlet country for each of the four conflict events, with term size reflecting relative frequency.
+These visualizations highlight the varied thematic focus across geopolitical actors.
+Figure 4: Thematic emphasis patterns across news outlets for four major conflict events, derived from GPT-3.5-tagged concepts, actors, and subtopics aggregated by outlet country.
+Temporal Sentiment Analysis:
+Figure 5 shows 5-day rolling average sentiment scores toward Russia and Ukraine in the immediate aftermath of the Bucha massacre (Mar - Apr 2022), aggregated by source country.
+Ukrainian media exhibits consistently negative sentiment toward Russia, as expected given their direct involvement in the conflict.
+Occasional positive spikes likely reflect optimistic coverage of military resistance or international support pledges.
+Russian media also shows a deeply negative sentiment – counterintuitive, given state alignment.
+This pattern suggests that war reporting employs vocabulary (violence, casualties, destruction) that sentiment models flag as inherently negative, regardless of political framing.
+Media in the USA and the UK exhibit pronounced volatility, with positive sentiment spikes corresponding to perceived diplomatic wins such as the US President’s Warsaw speech or high-level visits to Kyiv.
+Chinese media maintains characteristically moderate sentiment, consistent with their “strategic partnership” framing, though scores decline during peak media coverage of Bucha in early April.
+These patterns illustrate how sentiment analysis captures both explicit editorial stance and the underlying semantic content of war reporting, which may skew negative even if events are framed favorably.
+Figure 5: The average sentiment scores at 5-day intervals, for Russia (left) and Ukraine (right), during the aftermath of the Bucha massacre (March - April 2022).
+News outlets are grouped by their national affiliations.
+Appendix C. Annotation Guidelines
+We share the guidelines provided to human annotators for the validation experiments (see § 4.4).
+Topical Framing Annotation: The task is to identify the topical frames expressed in an article using the full text as context.
+Articles may receive multiple labels reflecting co-occurring themes.
+The labels are defined as per Table 3.
+Annotators are instructed as follows:
+  • assign labels based on explicit or implicit framing cues in the text;
+  • multiple labels are permitted and encouraged when themes co-occur;
+  • provide confidence scores [0, 1] for each assigned label; and
+  • avoid ranking labels by importance.
+Stance Annotation: The task is to determine the stance expressed in a sentence toward a specified target, using the surrounding text as context. Each sentence receives exactly one of these labels:
   • (1)
-    direct statements of observable or confirmed events (factual);
+    favor: positive stance (support, praise, endorsement);
   • (2)
-    claims attributed to specific actors (attributed);
+    against: negative stance (criticism, blame, opposition); or
   • (3)
-    explicit denials or rebuttals (disputed);
-  • (4)
-    indirect accounts from third-party media (reported);
-  • (5)
-    official statements of intent or directive (command/policy); and
-  • (6)
-    interpretive or evaluative judgments by experts or institutions (analytical/assessment).
-With GPT-3.5 plus human-in-the-loop validation, we label claims via structured prompts, and group semantically related claims through agglomerative clustering on Sentence Transformer representations Reimers and Gurevych (2019).
-This approach reveals distinct narrative patterns across geopolitical boundaries: Russian sources might characterize an airstrike as precision targeting of military assets (factual/attributed), while American, British, or Ukrainian outlets describe the same event as civilian infrastructure targeting, often citing eyewitnesses (disputed/reported).
-Analyzing claim pairs shows entailment occurring primarily within aligned media ecosystems: Ukrainian and Western sources reinforce each other through overlapping factual, reported, or command/policy claims, while Russian media maintains internal consistency through attributed and analytical claims citing official statements.
-Contradictions predominantly emerge across geopolitical divides as competing interpretations rather than direct disputes.
-Outlets rarely offer explicit rebuttals, instead employing different topical framing, omission, or counter-attribution strategies.
-Table 7 presents contradictory claim pairs from Russian and Ukrainian media.
-While referring to the same events, the claims construct incompatible interpretations of responsibility and action through attribution and selective emphasis.
-6 Conclusion
-The dnipro corpus represents a significant advancement in resources for computational analysis of media narratives during international conflicts.
-By integrating diverse geopolitical perspectives, consistent translations, and structured metadata across 246K articles, this dataset enables systematic study of narrative divergence that previous resources could not support.
-Our exploratory applications demonstrate dnipro’s utility across multiple analytical dimensions, from stance and sentiment detection to topical framing and contradiction analysis.
-The findings reveal how global media ecosystems construct parallel realities through attribution choices, selective emphasis, and strategic framing, rarely engaging with competing narratives.
-Beyond its utility in computational journalism, dnipro provides a foundation for developing more sophisticated methods to detect cross-cultural narrative divergence, identify emergent propaganda strategies, and understand information warfare dynamics.
-As the global information landscape grows increasingly complex, such resources become essential for technical advancement as well as a broader understanding of narratives that shape our perception of international conflicts.
-limitations and ethics don’t count towards the page limit
-7 Limitations
-The Russo-Ukrainian War has not concluded.
-The dataset, which currently spans from Feb 2022 to August 2024, may omit important narratives and developments resulting from politically significant events thereafter (e.g., foreign policy changes of a powerful geopolitical entity).
-For the chosen news sources, a wider variety of publications representing NATO nations would be beneficial.
-Currently, only two of the thirty-two NATO nations are represented (the United Kingdom and the United States of America).
-There are other media sources like DW in Germany, which are highly referenced globally, so they are an influential factor.
-Additionally, representation of more countries with ambiguous involvement in the war would add value to a truly global version of the dataset.
-Potentially interesting countries would be Hungary and Romania, both of which border Ukraine; Belarus, also bordering Ukraine, but known to be an ally of the Russian state; Turkiye, a powerful regional actor, especially in the Black Sea; and Georgia, a NATO contender that continues to maintain reasonably strong relations with Russia.
-Another facet limiting dnipro is the imperfect ground-truth annotations for stance, and other discursive phenomena.
-While this is not a limitation for qualitative analyses or mixed-method approaches, it does hinder nuanced supervised learning directly from the corpus.
-Our exploratory analyses, while illustrative, are not exhaustive and serve primarily to demonstrate the dataset’s potential rather than to make definitive empirical claims.
-8 Ethical Considerations
-dnipro contains potentially sensitive content related to an ongoing conflict with intense geopolitical implications. We acknowledge several potential ethical dimensions:
-  • Data collection: All content was gathered from publicly available news sources without breaching access restrictions or terms of service.
-We do not include any copyright-protected content or any private communications.
-  • Bias, representation, and inclusion: Our corpus inherently reflects the biases present in global media coverage.
-The inclusion of opposing narratives—including state-controlled media—is a deliberate choice, to empower comparative analysis.
-We strongly caution users against interpreting the mere inclusion of a source as endorsement of its content.
-  • Potential misuse: While designed for media analysis, computational linguistics, and conflict studies, this resource could also be used for propaganda analysis.
-And thus, it could also be misused for propaganda analysis.
-We strongly urge the community to utilize this corpus for constructive research into understanding narrative differences to bridge societal gaps, rather than weaponize any findings.
-  • Lost in translation: Machine translation introduces potential distortions, particularly with politically charged terminology and diplomatic language use.
-We document the details of our translation process.
-Further, we retain the original articles for enhanced scrutiny and possible resolution by experts in Russian, Ukrainian, and Mandarin Chinese.
-  • Contextualization: We provide robust metadata (publication date, source country, outlet ownership) to ensure articles can be interpreted within their proper contexts rather than in isolation.
-By openly addressing these considerations, we aim to enable valuable research on contentious discourse while minimizing potential harms.
-It is our expectation that through the study of such discourse, researchers will be able to bring empirical enhancements to conflict resolution mechanisms.
-Applications of such datasets are continuously evolving, as is the dataset itself.
-As such, we earnestly welcome community feedback on additional safeguards.
-Appendix A Appendix
-A.1 Prompt Design Details
-For our quantitative analysis, we used the mistral-small-latest model via the Mistral API to jointly extract stance targets and classify stances, and employed the siebert/sentiment-roberta-large-english model from Hugging Face to estimate sentiment polarity toward named entities.
-We used GPT-3.5 (via the OpenAI API) to assist with annotation for three tasks: article-level tag extraction, claim identification, and claim type classification.
-Prompt configurations varied by task.
-We employed both zero-shot and few-shot strategies, along with system-level instructions and output formatting constraints, following best practices in prompt engineering for language models Saravia (2022).
-A.1.1 Annotation for Stance
-For stance detection, we used the mistral-small-latest model via the Mistral API.
-The task was to classify whether a piece of news content expressed a stance FAVOR, AGAINST, or NONE toward a generated target.
-We followed a combined target generation at first followed by stance classification.
-Akash et al.
-(2025) We ran these experiments on 501 articles that mentioned the Bucha massacre dating from March to April 2022.
-The following zero-shot prompt was used to guide the model:
-Prompt:
-Listing 1: Zero-shot prompt for joint target generation and stance classification.
-Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a specified target.
-Analyze the following text, generate the target for this text, and determine its stance toward the generated target.
-A target should be the topic discussed in the text.
-The target can be a single word or a phrase, but its maximum length must be 4 words.
-If the stance is in favor of the generated target, write FAVOR; if it is against the target, write AGAINST; if it is ambiguous, write NONE.
-The answer must be exactly one of these three words: FAVOR, AGAINST, or NONE.
-Output format:
-Target: <target>, Stance: <stance>
-A.1.2 Annotation for Sentiment
-We used the siebert/sentiment-roberta-large-english model from Hugging Face to perform sentence-level sentiment classification on all news articles.
-For each entity of interest (e.g., Russia, Ukraine, Bucha), we first extracted all sentences from the article that mentioned the entity, using spaCy’s English language model for sentence segmentation.
-Prompt:
-What is the sentiment(positive, negative, or neutral) towards {entity} in this context: ’{text}’?
-A.1.3 Annotation for Article-level Tags
-To annotate articles with high-level thematic tags, we used a zero-shot prompt with GPT-3.5 that instructed the model to extract a set of concise, specific tags summarizing each article’s content.
-Tags were intended to capture relevant people, locations, military units, political themes, and ideological framings.
-To ensure consistency and support post-processing, the prompt enforced a structured JSON output format.
-Prompt:
-You are a geopolitical analyst trained to extract concise topical tags from news articles related to international conflict, especially the Russia–Ukraine war.
-Your task is to read the content of a single article and extract 5 to 12 relevant tags that capture the article’s key topics, actors, places, events, or narrative themes.
-Each tag should be 1–4 words long, lowercase unless a proper noun, and specific rather than general.
-Avoid vague tags like “fake”, “news”, or “war” unless central to the article.
-Tags may include named individuals, military groups, operations, locations, events, or ideological narratives.
-Output the result in valid JSON using the following structure:
-  "article_123": [
-    "bucha incident",
-    "sergei lavrov",
-    "false flag allegations",
-    "media manipulation",
-    "un security council"
-A.1.4 Annotation for Claim Extraction with Summary and Claim Types
-We used a structured few-shot prompt with GPT-3.5 to extract and classify claims from news articles.
-Each prompt instructed the model to identify distinct claims within an article, summarize them neutrally, and assign a claim_type based on a predefined taxonomy.
-Claims could span one or more sentences and were returned in a structured JSON format to support post-processing.
-Prompt:
-You are a geopolitical analyst trained to extract claims from news articles about the Russia–Ukraine war.
-Your task is to read the content of a single article and extract each distinct claim it makes.
-For each claim, return the original sentence(s), a short neutral summary, and a claim_type.
-Claims may span multiple sentences.
-Use only the article content field for analysis.
-The claim_type must be one of the following: Factual, Attributed, Disputed, Reported, Command/Policy, or Analytical/Assessment.
-Output the result in valid JSON using the following structure:
-  "article_123": {
-    "claim_1": {
-      "sentence":
-        "Ukrainian Side Reports Multiple
-        Casualties in Kyiv Airstrike.",
-      "summary":
-        "Ukraine reports casualties in
-        Kyiv airstrike.",
-      "claim_type": "Attributed"
-    "claim_2": {
-      "sentence":
-        "Zelensky Orders: ’Ensure NATOs
-        Long-Term Presence in the Black
-        Sea’.",
-      "summary":
-        "Zelensky orders long-term NATO
-        presence in the Black Sea.",
-      "claim_type": "Command/Policy"
-A.2 Use Case Experiment Details
-Figures 4(a) and 4(b) illustrate 5-day interval favor stance percentages toward Russia and Ukraine across news source countries during March–April 2022.
-(a) favor stance towards Russia (by source country).
-(b) favor stance towards Ukraine (by source country).
-Figure 4: 5-day interval favor stance percentages toward Russia and Ukraine (during the Bucha massacre, March–April 2022). News outlets are grouped by their national affiliations.
+    none: No clear stance, or purely descriptive text, or ambiguous stance.
+The annotators are provided these instructions:
+  • Use the full passage to resolve ambiguous pronouns or implicit references.
+  • If the target is none, the stance is also none.
+  • Focus on the stance expressed in the sentence, not your personal opinion.
+  • Implicit stances (e.g., through selective vocabulary) should be labeled.
+Table 8 and Table 9 provide annotated examples from the gold-standard human-annotated validation set, illustrating the topical framing and stance annotation tasks, respectively.
+For stance annotation, each sentence is paired with a stance label assigned toward a target entity.
+For topical framing, each passage may be associated with one or more topical frames, reflecting the co-occurrence of themes within the same article.
+The highlighted spans indicate the textual evidence that motivated the assigned labels.
+These highlights are intended to make the annotation rationale explicit and to ease reader interpretation; they were not provided to human annotators, and as such, the annotations are not derived from any span-level supervision.
+Passage Frames
+Title: Putin gives honorary title to Russian brigade accused of war crimes in Bucha
+Excerpt: A Russian brigade accused of committing war crimes in Bucha was awarded an honorary title by President Vladimir Putin.
+Putin congratulated the unit for its "great heroism and courage" and awarded it the title of "Guards" for "protecting Russia’s sovereignty." International leaders condemned the alleged atrocities, while investigators cited "reasonable grounds" to believe war crimes were committed.
+$\bullet$ morality: 0.9
+$\bullet$
+politics: 0.6
+$\bullet$
+legal: 0.6
+$\bullet$
+security: 0.6
+Table 8: Topical framing annotation examples showing multi-label assignments with confidence scores.
+Sentence (with target) Stance
+Russia has failed to capture Kyiv or Kharkiv, Ukraine’s two biggest cities, and suffered heavy losses against a far stronger Ukraine than even its allies expected. favor
+Aurel Braun, a professor at the University of Toronto who has written presciently on how western governments and companies misjudged the irrationalities of the Putin regime, points out that international relations is full of theories about how the interplay of powerful corporate interests and government produces policy.
+against
+When we see Russia’s unprovoked and unjustified invasion of Ukraine and massive disinformation campaigns and information manipulation, it is essential to separate lies from facts. against
+US officials have calculated that Putin, facing crippling sanctions across other sectors of his economy, can’t afford to weaponize his energy supplies by cutting off sales to Europe. none
+Vladimir Putin awarded Heroes of the Russian Federation in the Yekaterininsky Hall. none
+Table 9: Stance annotations from the gold-standard human validation set, shown with highlighted target entities.